### v1

# Gianluca Monaci          Rafael S. Rezende          Romain Deffayet          Gabriela Csurka Guillaume Bono          Hervé Déjean          Stéphane Clinchant          Christian Wolf

Abstract
Methods for navigation based on large-scale learning typically treat each episode as a new problem, where the agent is spawned with a clean memory in an unknown environment. While these generalization capabilities to an unknown environment are extremely important, we claim that, in a realistic setting, an agent should have the capacity of exploiting information collected during earlier robot operations. We address this by introducing a new retrieval-augmented agent, trained with RL, capable of querying a database collected from previous episodes in the same environment and learning how to integrate this additional context information. We introduce a unique agent architecture for the general navigation task, evaluated on ObjectNav, ImageNav and Instance-ImageNav. Our retrieval and context encoding methods are data-driven and heavily employ vision foundation models (FM) for both semantic and geometric understanding. We propose new benchmarks for these settings and we show that retrieval allows zero-shot transfer across tasks and environments while significantly improving performance.
1 Introduction
A realistic setting for robot navigation is the “unboxing” scenario, where one robot (or a fleet) is placed in its target setting, started-up, and navigates “out of the box”. The main appeal of this scenario is the lack of dependency on environment preparation, like scanning the environment, installing external localization systems or generating a floor plan prior to navigation. Most modern methods tackling this scenario, e.g. [75, 12, 79, 6], use learning-based approaches in an episodic setting: every episode in the robot’s life is treated as if it was its first. This is obviously sub-optimal: during operation the scene is explored, objects and their locations are observed, affordances found, failure cases encountered etc. Exploiting this information is crucial for designing efficient navigation agents that can continuously improve.
Figure 1: We augment a navigation agent with context information retrieved from a scene-specific retrieval database, fed with data by one or potentially multiple robots. We leverage semantic and multi-modal foundation models for retrieval, and geometric foundation models for context encoding, harnessing its power to provide directional information. We explore ObjectNav, ImageNav and Instance-ImageNav tasks, including zero-shot cases.
Benchmarks like the k-item scenario [4], multi-object navigation [73] or GOAT benchmark [39] test the capacity of an agent to retain information from previous (sub-)goals and trajectories. Structured latent memory trained end-to-end with RL performs best in these benchmarks [50]. However, these architectures have only been tested on short (multi-)episodes of $500$ to $2,500$ agent steps and are hardly suitable for a real continuous operation: transformer-based agents with self-attention over time suffer from limited context length and the quadratic complexity of attention, while recurrent models are limited by the size of their latent memory and the network capacity growing quadratically with representation size. Alternative ways to integrate information from initial rollouts are topological maps [63, 69], which require the estimation and generation of a structured graph model from a sequence of posed observations. The extension of these methods to multi-robot scenarios is anything but trivial.
In this work we propose a simple data-driven approach to extend a state-of-the-art navigation agent [6] working out-of-the-box, with capabilities to store visual observations in a global indexed database, share them with a fleet, and retrieve and process them to improve performance during operation. The proposed architecture is general and can be used for different navigation tasks — we evaluate our work on ObjectNav, ImageNav and Instance-ImageNav. While the goal may be specified either as an image or a semantic category, depending on the task, the retrieved information is always visual first person views (FPVs). Computer vision is at the heart of the agent in multiple ways: data-driven retrieval mechanisms leverage vision foundation models (FM) like CLIP and DINOv2 to retrieve candidate views, while a geometric FM [74] integrates the retrieved context and provides directional information to the agent, cf. Fig. 1.
This approach is scalable, as the dataset can potentially be very large and can be queried by any number of agents with sub-linear time algorithms. Retrieving from an image database is also considerably simpler and more flexible than using involved memory structures such as metric or topological maps. The latter require pose to register observations and synchronization to enable multi-agent data collection: unless pose estimates are absolute (which requires scene preparation), fusing new updates from multiple agents requires a jointly maintained map and distributed algorithms.
We propose the following contributions: (i) new navigation tasks, where the environment is augmented with a retrieval database fed by one or multiple robots; (ii) a unified end-to-end architecture for Retrieval-Augmented Navigation, RANa, for multiple tasks; (iii) data-driven retrieval mechanisms that do not require any meta-data beyond raw FPVs; (iv) mechanisms that leverage retrieved information either for additional context, or to replace the goal on the fly, allowing for zero-shot applications, cf. Fig. 2.
Figure 2: We target several contributions: perf. improvement (Success +x), opening new use cases from existing models, and zero-shot applications with retrieval only added at test time (Refer to caption).
Figure 3: We propose an architecture for Retrieval-Augmented Navigation architecture – RANa, that addresses all use cases of Fig. 2. Dashed arrows are optional depending on the navigation task.
2 Related Work
Visual navigation – The task of navigation has been addressed in robotics research using mapping and planning [71, 46], which requires solutions for mapping and localization [9, 44, 54], planning [41, 66] and low-level control [26, 62]. These methods depend on accurate sensor models, filtering, dynamical models and optimization schemes. In contrast, end-to-end models learn deep representations such as flat recurrent states [79, 6], occupancy maps [13], semantic maps [12], latent metric maps [56, 4, 32, 50], topological maps [3, 14, 67], self-attention [16, 23, 25, 61] or implicit representations [51]. Then, they map these representations into actions, using RL [36, 53, 8], imitation learning [21] or by maximizing navigability [7].
In this work, our goal is to address end-to-end learning of representations and policies with RL. Although this problem is nearly solved for PointNav, where the agent must navigate to a relative GPS coordinate [64, 75], it remains a challenge for other goal specifications. ObjectNav, where the goal is an object category, requires semantic understanding of the scene, therefore a large body of research builds a semantic map [27, 12] or leverages an LLM to guide navigation based on a scene’s common sense semantics [83, 10, 43, 82]. Several approaches address this task in a zero-shot fashion, training their agents without ObjectNav rewards and leveraging CLIP encoders for goal representation [70, 47, 27]. Recently, Zeng et al. [84] tackles ObjectNav with a transformer state encoder on simulated environments.
In ImageNav, an image is provided as goal, requiring a geometric understanding of the scene. A common variant is Instance-ImageNav [42], where the goal image depicts a specific object in the environment and is taken with camera parameters different from those of the agent. Recently, the end-to-end model DEBiT [6] achieved state-of-the-art performance in ImageNav leveraging the geometric FM CroCo [74] to estimate relative poses between goal and agent’s observation. In this work, we build upon DEBiT by both extending it to ObjectNav and augmenting it with retrieval, naturally extending it to the multi-robot setting and enabling zero-shot ObjectNav and Instance-ImageNav.
To the best of our knowledge no other method performs all these tasks competently, the closest competitor being [47] which tackles ImageNav and zero-shot ObjectNav, although with considerably lower performance (see Sec. 5).
Method and Img Obj Not req.: Easiness of Multiple
representation Nav Nav pose depth repr. update collectors
GOAT [39] $\sim$ map reg.
3D Scene graphs [82] $\sim$ graph+map reg.
Topo map [63, 14, 67] $\sim$ graph upd.
Semantic maps [12] $\sim$ map reg.
Sem. Implicit [51] $\nabla$ upd
Feature SLAM [54] $\sim$ pose graph
Dense SLAM [24] $\sim$ map reg.
RANa store
To support multiple collectors, a method is either pose free, like RANa, or uses distributed algorithms to synchronize, or absolute pose (and not episodic).
Table 1: RANa compared to SotA in continual navigation. RANa is a flexible, data-driven solution that can easily exploit data from previous episodes, in particular collected by multiple robots.
Retrieval-augmented control – The idea of retrieving information from a dataset of past experiences has been introduced in earlier works. Among them, episodic control methods for RL [5, 57, 45, 29, 30, 33] act on successful experiences by re-employing Q-value estimates, enhancing sample efficiency during training. In contrast, we exploit retrieved information from past experiences of a robot fleet to enhance navigation performance and to enable zero-shot task generalization. Goyal et al. [28] augment an RL agent with a retrieval process parameterized as a neural network that has access to a dataset of past trajectories. Here instead we take inspiration from [34] that, to our knowledge, was the first work to propose a generic non-parameterized retrieval augmentation for an RL agent with previously learned embeddings. However, alleviating the need to pre-train embeddings using RL, we use vision foundation models to process stored and retrieved images. This enables zero-shot transfer and allows to adapt the approach to various navigation tasks.
Retrieval-Augmented Planning [38] uses a memory of successful task executions, comprising plans, actions and observation sequences. Past experiences relevant to the task at hand are leveraged to guide planning with an LLM agent based on task constraints and state similarity. Xie et al. [78] build a hierarchical topological memory to prompt an LLM for reasoning and goal-finding in a navigable environment. In contrast, we only store images derived from unstructured observations without any associated metadata.
Navigation supported by prior rollouts – Other navigation methods exploit previous observations to construct a topological map of the explored scene, and use it to guide navigation [63, 14, 67]. In contrast, our database is unstructured and we can query it to retrieve goal and context images with simple image retrieval.
In Tab. 1 we compare RANa with various navigation methods discussed above in the continual navigation scenario. Most approaches require depth and pose and/or more involved (metric or topological) map updates, which need at least to register observations. In contrast, RANa does not require depth or pose, the database update and retrieval is very simple, and naturally supports multiple data collection robots without the need of synchronization. While SLAM-based solutions [24, 54] are well established for continual navigation, with RANa we propose a lightweight data-driven solution that can be easily scaled up to take inputs from multiple robots and can tackle ImageNav, Instance-ImageNav and ObjectNav use cases.
3 Retrieval-augmented agent
We target navigation in 3D environments, where an agent is tasked to navigate from a starting location to a goal and receives at each timestep $t$ an image observation $\mathbf{x}_{t}$. Our method is general and the agent can be trained for a diverse range of tasks, but without loss of generality in this work we focus on ImageNav, where the agent receives an additional image goal and ObjectNav, where the agent receives an object goal which is a category in form of text, cf. Fig. 2. In both cases we denote the original goal by $\mathbf{g}$.
The action space is discrete, $\mathcal{A}$ ={MOVE FORWARD 0.25m, TURN LEFT $10^{\circ}$, TURN RIGHT $10^{\circ}$, and STOP}. Navigation is considered successful if the STOP action is selected when the agent is within 1m of the goal position in terms of geodesic distance. ObjectNav has two additional actions, LOOK UP and LOOK DOWN.
Our tasks are based on the ImageNav and ObjectNav tasks of the Habitat simulator and platform [64]. We extended them with a retrieval dataset $\mathcal{D}$ and a retrieval mechanism $\mathcal{R}$, which will be detailed in Sec. 4. The dataset $\mathcal{D}=\{\mathbf{x}_{i}^{\mathcal{D}}\}$ contains FPVs indexed by $i$ and stemming from previous episodes or from exploratory rollouts in the scene. While it is in principle possible to store additional meta-data associated to data collection, like approximate pose, performed actions, reward, success or value (in an MDP sense, when available), we show in this work that the access to FPVs alone provides rich information to the agent (see Sec. 4).
The agent queries this dataset $\mathcal{D}$ with the retrieval mechanism $\mathcal{R}$, which can be potentially used at each time step $t$. $\mathcal{R}$ uses the goal $\mathbf{g}$, and optionally the current observation $\mathbf{x}_{t}$, to retrieve a set of FPVs $\mathcal{R}(\mathbf{x}_{t},\mathbf{g})\longrightarrow\{\mathbf{r}_{t,1},\mathbf{r}_{t,2},\dots\mathbf{r}_{t,N}\}$, that is used as context information to improve navigation. Depending on the use case, the use of observation $\mathbf{x}_{t}$ in the retrieval process may be optional — if not provided, retrieval can be done once per episode, as opposed to every time step.
[…]
$\begin{array}[]{lll}\tilde{\mathbf{x}}_{t}&=x(\mathbf{x}_{t})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// perceive}}\\ \tilde{\mathbf{g}}_{t}&=g(\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// compare obs+goal}}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{g}}_{t},l(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// update rec state}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// act}}\\ \end{array}$
where $x$ and $g$ are trainable encoders, $h$ is the update function of a GRU [18] which maintains a hidden state $\mathbf{h}_{t}$ over time; $l$ is an embedding function, and $\pi$ is the policy. For clarity we have omitted the equations of the gating functions.
Then, the retrieval-augmented agent is defined as:
$\begin{array}[]{lll}\tilde{\mathbf{x}}_{t}&=x(\mathbf{x}_{t})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// perceive}}\\ \tilde{\mathbf{g}}_{t}&=g(\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// compare obs+goal}}\\ \tilde{\mathbf{c}}_{t}&=c(\mathcal{R}(\mathbf{x}_{t},\mathbf{g}),\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// retrieve context}}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{g}}_{t},\tilde{\mathbf{c}}_{t},l(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// update rec state}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// act}}\\ \end{array}$
where $c$ is an encoder which compresses the retrieved context into a compact form. It is trained end-to-end together with the agent, as described in Sec. 3.1, while the retrieval process $\mathcal{R}$, described in Sec. 4, leverages pre-trained models. In all tasks, the observation encoder $x$ is implemented as a half-width ResNet-18 (hwRN18) architecture [31].
The goal $\mathbf{g}$, which is an image or a category, according to the navigation task, is compared to the observation $\mathbf{x}_{t}$ through a dedicated comparison function $g$ defined as
$g(\mathbf{x}_{t},\mathbf{g})=\left\{\begin{array}[]{ll}d(\mathbf{x}_{t},\mathbf{g})&{\scriptsize\texttt{if ImageNav}}\\ [cl_{V}(\mathbf{x}_{t}),cl_{T}(\mathbf{g})]&{\scriptsize\texttt{if ObjectNav}}\\ \end{array}\right.$ (3)
As in [6], for ImageNav we use a binocular visual encoder $d$ that compares the observation $\mathbf{x}_{t}$ with the goal image $\mathbf{g}$ and returns an embedding representing information about the goal’s direction, pose, and visibility. This encoder $d$ has been fine-tuned from the foundation model CroCo [74], solving as pre-text tasks relative-pose and visibility estimation [6]. We use the weights available online. For ObjectNav where a goal category is specified, we use a CLIP encoder [59] providing a semantic comparison of the observed image, encoded as $cl_{V}(\mathbf{x}_{t})$, and the object category, encoded as $cl_{T}(\mathbf{g})$.
Additionally, an image goal $\mathbf{g}^{\mathcal{R}}$ can be retrieved from $\mathcal{D}$ given the task goal $\mathbf{g}$, allowing the ImageNav model to address tasks in a zero-shot way. For example, an ImageNav agent can perform ObjectNav by retrieving a goal image $\mathbf{g}^{\mathcal{R}}$ from the database $\mathcal{D}$ given a goal category ([Uncaptioned image], [Uncaptioned image]). This can be integrated into the augmented agent above by adding a third case to the goal comparison function $g$,
$g(\mathbf{x}_{t},\mathbf{g})=d(\mathbf{x}_{t},\mathbf{g}^{\mathcal{R}})\leavevmode\nobreak\ \leavevmode\nobreak\ {\scriptsize\texttt{if image goal isretrieved}}$ (4)
We explore this use of retrieval for zero-shot task generalization in Sec. 5.2.
In summary, the agent receives at each timestep a new visual observation $\mathbf{x}_{t}$, which it encodes, compares with the goal $\mathbf{g}$, and enriches with additional context information retrieved from the database $\mathcal{D}$. This representation is then passed to a recurrent policy. See Fig. 3 for a visual overview.
3.1 Trainable context encoder $c$
[…]
$\mathbf{e}_{t,n}=\leavevmode\nobreak\ \leavevmode\nobreak\ d(\mathbf{x}_{t},\mathbf{r}_{t,n})$ // ObjNav (5)
$\mathbf{e}_{t,n}=[d(\mathbf{x}_{t},\mathbf{r}_{t,n})\ d(\mathbf{g},\mathbf{r}_{t,n})].$ // ObjNav or ImageNav (6)
At this point, we obtain a list of embeddings $\mathbf{e}_{t,n}$ that benefit from the rich inductive bias provided by the geometric FM, including information about pose and visibility of the context image $\mathbf{r}_{t,n}$ from the observed image $\mathbf{x}_{t}$ and goal image $\mathbf{g}$. We must now derive a single context embedding $\tilde{\mathbf{c}}_{t}=c(\mathcal{R}(\mathbf{x}_{t},\mathbf{g}),\mathbf{x}_{t},\mathbf{g})$ that provides useful information to the agent. We opt for two variants:
Variant 1: Gumbel soft-max selector – The agent selects a single embedding from the list by predicting a distribution over items in the retrieved list followed by a Gumbel soft-max sampler [37]. More precisely,
$\tilde{\mathbf{c}}_{t}\sim\sigma(\{\alpha_{n}\}_{n=1}^{N}),\quad\alpha_{n}=\texttt{Linear}(\mathbf{e}_{t,n})$ (7)
where $\sigma$ is the soft-max distribution over items in the list. We perform the sampling and the computation of gradients using the Gumbel soft-max trick [37]. We motivate this choice from the pre-training of the geometric FM $d$ in Eq. 5 and Eq. 6: the latent vector $z=d(a,b)$ has been pre-trained to provide overlap information between images $a$ and $b$, which we conjecture to be highly correlated with the relevance of the corresponding context item. Additionally, since the weights $\alpha_{i}$ are calculated independently, the model can be zero-shot applied to any context size at test time.
Variant 2: attention-based encoder – A transformer-based model is potentially capable of integrating multiple context tokens $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ into a single fused embedding, effectively changing and enriching the original embedding space. We use a transformer encoder network with 2 layers of 4-heads self-attention and predict the output by concatenating the context tokens and processing them with a 2-layer MLP.
4 Retrieval mechanisms for RANa
In all tasks and use cases, the retrieval process consists of a nearest neighbor search on representations obtained with pre-trained models. Depending on the navigation task at hand, we either use DINOv2 [55] when we have a visual query or Open-CLIP [35] for textual queries. For zero-shot use-cases that require replacing the task goal $\mathbf{g}$ with a retrieved image goal $\mathbf{g}^{\mathcal{R}}$ ([Uncaptioned image], [Uncaptioned image]), we set $\mathbf{g}^{\mathcal{R}}$ as the nearest neighbor of $\mathbf{g}$ from the dataset $\mathcal{D}$. For use-cases that require building a context set to be passed to the encoder $c$ ([Uncaptioned image], [Uncaptioned image], [Uncaptioned image]), we investigate two strategies: (i) static context, where only the goal is used to build a fixed context kept throughout the episode, and (ii) dynamic context, which also depends on the current observation and is recomputed at each timestep.
4.1 Static context
We build the static context by ranking images from the dataset based on DINOv2 or CLIP feature similarity with the goal and selecting the top-$N$ nearest neighbors.
Diversity – providing multiple near-duplicates to the selector might not be very helpful. Moreover, in ObjectNav a scene may contain several instances of the target class and it is better to provide context images showing different instances. We therefore resort to Maximum Margin Relevance (MMR) [11] re-ranking to increase diversity. Assuming a relevance score vector $\omega$ (DINOv2 or CLIP similarities) providing a ranked shortlist (top-100 in our experiments), and a matrix $\Omega$ encoding the similarity for each pair of images in the ranked shortlist, MMR greedily selects at each step (rank) $i$ the element $r_{i}$ that maximizes the re-ranking criterion:
${\text{MMR}(r_{i})=\beta\omega(r_{i})-\left(1-\beta\max_{r_{j}\in\mathcal{S}_{i-1}}\Omega(r_{i},r_{j})\right)},$
where $\beta\in[0,1]$ is a mixture parameter (set to $0.5$ in all our experiments) and $\mathcal{S}_{i-1}$ is the set of images already selected. We then select the top-$N$ highest MMR.
CLIP similarity normalization – While CLIP allows to easily compare any open-vocabulary object category with retrieval gallery images, it has been observed in the context of zero-shot image classification that directly ranking images by similarity to the category is suboptimal [58]. To address this issue for ObjectNav, we leverage the fact that we often target a pre-defined set of classes $\mathcal{O}$ (e.g. 6 in HM3D), and rescale the score of each image in $\mathcal{D}$ via softmax:
${\omega_{\mathbf{o}}(\mathbf{x})=e^{cl_{S}(\mathbf{o},\mathbf{x})}\,/\,{\left(\sum_{\mathbf{o}^{\prime}\in\mathcal{O}}e^{cl_{S}({\mathbf{o}^{\prime},\mathbf{x}})}\right)}}$
where $cl_{S}(\mathbf{o},\mathbf{x})$ denotes the CLIP similarity of image $\mathbf{x}$ and object category $\mathbf{o}$ and $\omega_{\mathbf{o}}(\mathbf{x})$ is its normalized similarity.
4.2 Dynamic context set
It can be interesting to build the context using not only the goal but also the current observation, for example to find intermediate waypoints between the current observation and the goal. We propose to use an affinity graph based on similarities of images in the dataset $\mathcal{D}$ to find such waypoints.
DINOv2-Graph – Given $\mathcal{D}$, we can build the affinity matrix $\Omega_{\mathcal{D}}$ containing the similarity between all pairs of images in $\mathcal{D}$ (we use DINOv2 [55] features). Using $\Omega_{\mathcal{D}}$, we can build a graph whose nodes are images $\mathbf{x}_{i}^{\mathcal{D}}\in\mathcal{D}$ and edges are weighted by similarity. Then, we find in $\mathcal{D}$ the most similar images $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ to the observation $\mathbf{x}_{t}$ and goal $\mathbf{g}$, respectively. Finally, we populate the context $\mathcal{R}(\mathbf{x}_{t},\mathbf{g})$ with $r_{1}^{\mathbf{x}_{t}}$ , $r_{1}^{\mathbf{g}}$ and images sampled on the shortest path in the graph between them. We provide more details about the graph and its relation to topological maps in the Supplementary.
5 Experimental results
We train and evaluate our agents on the Habitat simulator and platform [64] according to the standard ImageNav, Instance-ImageNav and ObjectNav task definitions.
Retrieval-augmented navigation benchmarks – There is no publicly available navigation benchmark that provides a database of observations that can be exploited during training and evaluation. We thus propose to derive ImageNav and ObjectNav benchmarks from standard Habitat tasks [64], modifying them as little as possible to enable training and evaluating retrieval-augmented agents. For ImageNav, we use the Gibson dataset [77], consisting of 72 train and 14 eval scenes, and for ObjectNav, we use 80 train and 20 eval scenes of the HM3DSem-v0.2 [60, 81]. All agents use only one $112{\times}112$ RGB sensor. For each scene, unless otherwise stated, we generate a retrieval set $\mathcal{D}$ of $1,000$ FPV images by letting agents navigate the environment and store images in $\mathcal{D}$. During training and evaluation, the retrieval-augmented agent can select dataset images from the scene it operates in. Further details on dataset creation are in Supplementary.
Implementation details – For ImageNav ([Uncaptioned image], [Uncaptioned image]), we use the experimental setup for Gibson dataset [68] where the goal is a random view taken by the agent camera, as in, e.g.[48, 52, 6]. We base our agents on the DEBiT architecture [6] by starting from the official codebase and weights provided by the authors [^2] https://github.com/naver/debit, and extend them as described in Sec. 3. We consider two DEBiT backbones, base (DEBiT-b) with a binocular encoder with 55M parameters, and tiny (DEBiT-t), with a 17M-parameters encoder to allow faster execution for the ablation studies on Sec. 5.4. The corresponding retrieval-augmented models are RANa-b and RANa-t. In Sec. 5.2 we use the same agent and weights in zero-shot fashion for Instance-ImageNav and ObjectNav when the task goal is replaced with a retrieved image goal ([Uncaptioned image], [Uncaptioned image]).
For the ObjectNav base architecture with a textual goal ([Uncaptioned image], [Uncaptioned image]), we use the default experimental setup of HM3DSem-v0.2 [60, 10]. We modify the RGB observations to be $112{\times}112$ instead of $480{\times}640$ as in the original ObjectNav definition, so we can load the pretrained weights of the hwRN18 visual encoder $x$ from DEBiT-t [6]. We use the OpenCLIP [35] model trained on Laion-2B [65] [^3] https://github.com/mlfoundations/open_clip/.
Model Retrieval Encoder SR SPL
[Uncaptioned image] DEBiT-t [6] 79.9 51.4
[Uncaptioned image] RANa-t DINOv2 Gumbel 83.0 52.3
[Uncaptioned image] RANa-t DINOv2 Attention 81.9 54.3
[Uncaptioned image] RANa-t DINOv2-Graph Gumbel 82.0 51.5
[Uncaptioned image] RANa-t DINOv2-Graph Attention 84.7 63.7
[Uncaptioned image] DEBiT-b [6] 83.4 56.8
[Uncaptioned image] RANa-b DINOv2 Gumbel 88.3 57.6
[Uncaptioned image] RANa-b DINOv2 Attention 84.3 58.3
[Uncaptioned image] RANa-b DINOv2-Graph Gumbel 89.7 60.4
[Uncaptioned image] RANa-b DINOv2-Graph Attention 90.7 71.8
[Uncaptioned image] hwRN18+CLIP 42.9 23.0
[Uncaptioned image] RANa-t OpenCLIP Gumbel 52.1 26.9
[Uncaptioned image] RANa-t OpenCLIP Attention 49.7 26.0
Table 2: Effect of retrieving context on ImageNav and ObjectNav. models finetuned for the same number of steps as corresponding RANa models for fairness of comparison.
Model training – All models, retrieval-augmented or not, are trained with PPO up to 200M steps. Unless specified otherwise, the geometric FM $d$ and encoders $x$ and $l$ are loaded from DEBiT and kept frozen, as are OpenCLIP and DINOv2. For retrieval-augmented agents, the context encoder $c$ and policy network $\pi$ are learned from scratch. Compared to DEBiT and hwRN18+CLIP , the GRU memory of RANa receives the additional context input $\tilde{\mathbf{c}}_{t}$, therefore we expand the DEBiT/hwRN18+CLIP GRU input matrix to reach the required dimension and finetune it. Further details and ablations of this choice are reported in Supplementary. The reward definition is inspired by PointGoal [15] and ImageNav [6] rewards and given as $r_{t}=\mathrm{K}\cdot\mathbf{1}_{\text{success}}-\Delta_{t}^{\mathrm{Geo}}-\lambda,$ where $K{=}10$, $\Delta_{t}^{\mathrm{Geo}}$ is the increase in geodesic distance to the goal, and slack cost $\lambda{=}0.01$ encourages efficiency.
Retrieval – We consider two context encoding approaches, Gumbel and Attention, and three retrieval mechanisms based on DINOv2 and DINOv2-graph for ImageNav, and OpenCLIP for ObjectNav tasks. Unless otherwise specified, we use a context size $N{=}8$. Results for ObjectNav use a retrieval dataset of size $10$k images (see Sec. 5.4 and Supplementary for ablations).
Metrics – Navigation performance is evaluated by the success rate (SR), i.e., the percentage of episodes terminated within a distance of ${<}1$m from the goal after the agent has called the STOP action, and SPL [2], i.e., SR weighted by the optimality of the path: $\textit{SPL}=\frac{100}{N}\sum_{i=1}^{N}S_{i}\frac{\ell_{i}^{*}}{\max(\ell_{i},\ell_{i}^{*})},$ where $S_{i}$ is a binary success indicator in episode $i$, $\ell_{i}$ is the agent path length and $\ell_{i}^{*}$ the shortest path length.
5.1 Retrieval improves navigation
Tab. 2 shows results of the proposed RANa agents with retrieved context on ImageNav and ObjectNav tasks, which exhibit significant performance improvements across all settings compared to the DEBiT baseline and the ObjectNav baseline models [^4] To make baselines comparable, the loaded checkpoints continued training for the same additional 200M steps updating GRU and policy only.. For static context retrieval (both using DINOv2 for ImageNav or OpenCLIP for ObjectNav) the Gumbel encoder shows better results, arguably because there is no inherent ordering in the context and the retrieved CroCo-based features contain strong information about the relative pose between context elements and observation. The attention-based encoder works better for the dynamic graph-based context, where reasoning about the order of context elements becomes important. The graph-based context combined with attention obtains the best results, with a huge boost in SPL ($+14.2$ and $+12.3$ points) compared to the DEBiT-b and DEBiT-t baselines, respectively. For all ObjectNav runs, we found that finetuning $x$ while training on HM3D improved performance, as the visual encoder of DEBiT was only trained on Gibson observations. Moreover, since ImageNav and ObjectNav have different action spaces, we learn the action embedding $l$ from scratch.
Model SR SPL Binocular encoder
ZSON [47] 36.9 28.0 Observation finetuned, goal frozen
OVRL [80] 54.2 27.0 Finetuned
VC1-ViT-L [49] 81.6 - Finetuned
OVRL-v2 [79] 82.0 58.7 Finetuned
DEBiT-b no adapters [6] 83.0 55.6 Frozen
DEBiT-L no adapters [6] 82.0 59.6 Frozen
DEBiT-L+adapters[6] 94.0 71.7 Finetuned with adapters
RANa-b Graph-Attn 90.7 71.8 Frozen (base DEBiT-b no adapters)
Table 3: ImageNav state-of-the-art results. these models use additional GPS+Compass input providing exact localization.
Comparison to ImageNav SotA – In Tab. 3 we compare RANa ImageNav with state-of-the-art (SotA) models that typically finetune (with a navigation loss) visual encoders based on CLIP [47], ViT [80, 49, 79] and geometric FM [6] used for observation-goal comparison. RANa boosts the performance of DEBiT-b, which uses frozen binocular encoders, and achieves performance on par with SotA method DEBiT-Large [6] finetuned with adapters [17], which use an involving adaptation process to tune visual encoders to the task. In contrast, we achieve strong performance through (i) scalable context encoding and (ii) a generic frozen visual FM, enabling several zero-shot tasks (Sec. 5.2).
Comparison to ObjectNav SotATab. 4 compares the results of RANa ObjectNav tiny model with competing approaches. Compared to methods of similar sensor settings (one FPV RGB camera), RANa-t outperforms previous SotA by $+9.7$ on SR and $+7.7$ on SPL. Neither ZSON [47] nor PSL [70] were trained on ObjectNav episodes, but on HM3D ImageNav episodes, which might give a small advantage to RANa. Hence we also compare RANa-t to methods that use extra sensors, such as depth camera, GPS+compass, or 6 camera sensors, as well as LLM-guided navigation. Here RANa-t shows competitive results, outperforming several strong methods such as ESC [86] and PixNav [10]. Compared to the considerably more complex L3MVN [83], which leverages depth and pose to construct a semantic map and uses GPT-2 to guide exploration via object searching, RANa-t is competitive, being $-2.1$ below on SR but $+1.4$ above on SPL, meaning that the retrieved context guides our agent to more efficient routes to object goals.
100 dataset images suffice to improve performanceTab. 5 shows that the size of the retrieval database can be small, $100$ is enough for the considered Gibson scenes in ImageNav. Additional experiments for ObjectNav and zero-shot tasks are presented in Supplementary.
Model Extra sensors w/ map w/ LLM SR SPL
PixNav [10] Pano RGB 37.9 20.5
ESC [86] Depth, odom 39.2 22.3
L3MVN [83] Depth, odom 54.2 25.5
CoW [27] Depth, odom 6.1 3.9
GoW [27] Depth, odom 32.0 18.1
ProcTHOR [19] Depth, odom 13.2 7.7
ZSON [47] 25.5 12.6
PSL [70] 42.4 19.2
RANa-t (Ours) 52.1 26.9
Table 4: Comparison to ObjectNav state-of-the-art methods.
100 1k 10k 20k
1k 81.7 (50.1) 82.7 (52.6) 82.1 (51.2) 82.8 (51.9)
10k 80.6 (49.4) 83.0 (52.3) 79.8 (46.3) 82.2 (51.5)
Table 5: [Uncaptioned image]  Impact of retrieval database size - SR (SPL) for ImageNav with retrieval using different database sizes, combinations of train and test conditions.
5.2 Retrieval allows zero-shot generalization
By retrieving the image goal $\mathbf{g}^{\mathcal{R}}$ from the database, it is possible to apply any ImageNav model to the ObjectNav and Instance-ImageNav tasks. Tab. 6 compares the performance of zero-shot ImageNav architectures in: (i) Gibson ObjectNav, a variant introduced in [1] for ImageNav agents, consisting of $1,000$ episodes over $5$ Gibson scenes and containing the 6 object categories of HM3D ObjectNav, and (ii) HM3D Instance-ImageNav with modified camera intrinsics and robot size to fit the ImageNav configuration. In both cases we use retrieval databases of $50,000$ images, since it is important to retrieve an image goal $\mathbf{g}^{\mathcal{R}}$ as close as possible to the goal, as shown in Sec. 5.4. In Instance-ImageNav we replace the image goal with the closest image in DINOv2 feature space from the retrieval set. In zero-shot ObjectNav we retrieve the top-9 closest images in CLIP space and re-rank them at each step to select as goal $\mathbf{g}^{\mathcal{R}}$ the element with the highest co-visibility with the current observation, and the remaining 8 images constitute the context. This is done using the frozen co-visibility auxiliary loss head used to train DEBiT models. This step is needed since in ObjectNav any instance of a target object is a valid goal.
Retrieval enables the use of DEBiT-b for both tasks, achieving strong performance in line with best existing methods (for reference, DEBiT-b achieves $10$ SR and $3.1$ SPL on the original Instance-ImageNav task). RANa outperforms the feature-matching based method [42] which uses additional depth and odometry sensors and is the current SotA in zero-shot Instance-ImageNav, as well as ZSON (configuration A) [47] in zero-shot Gibson ObjectNav.
ObjNav IIN
Model Extra sensors SR SPL SR SPL
ZER [1] 11.3
ZSON A [47] 31.3 12.0 16.1 8.4
Mod-INN [42] Depth, odom 56.1 23.3
[Uncaptioned image] DEBiT-b 31.9 15.8 55.4 25.5
[Uncaptioned image] RANa-b 36.6 17.9 58.7 26.8
Table 6: Zero-shot navigation results using retrieved image goal $\mathbf{g}^{\mathcal{R}}$ for Gibson ObjectNav and HM3D Instance-ImageNav, and comparison to the state of the art.  obtained with camera intrinsics and robot size of the original Instance-ImageNav task specification.
Fig. 4 shows an example of successful zero-shot Instance-ImageNavigation, where our RANa agent first correctly picks from the context the door near the goal ①, and then a view of the room where the target is ②, leading to success. In contrast, the DEBiT-b baseline without context does not see the target inside the room and passes by, exploring the wrong side of the house and failing to complete the episode.
Figure 4: Zero-shot Instance-ImageNavigation episode by RANa and by the DEBiT-b baseline without context (top-right box). The green camera icon indicates the pose of the retrieved image goal $\mathbf{g}^{\mathcal{R}}$, the pink ones indicate the poses of selected context items.
5.3 Retrieval allows bridging the sim-to-real gap
We evaluate our [Uncaptioned image] RANa-b agent on a small Instance-ImageNav dataset of $200$ episodes where goal images are pictures taken in a real office environment. We manually annotate the positions of $20$ real images in a simulated 3D reconstruction of the environment. We generate a retrieval set $\mathcal{D}$ of $10$k images from uniformly-sampled navigable poses, rendered in simulation using the agent configuration. Simulating navigation from $10$ random starting poses towards each of the $20$ real image goals, RANa reaches $46.5$ success, vs. $40.5$ of DEBiT-b (see inlet Table in Fig. 5). The Figure illustrates a challenging success case (left), where the cactus viewing angle is far from any view the agent could capture, and (right) an interesting failure case, where the retrieval returned as goal $\mathbf{g}^{\mathcal{R}}$ an image of a different instance of the same microwave model at a different position. Interestingly, some retrieved context images depict the correct goal, and RANa appears to be able to exploit this information as it succeeds in $3/10$ episodes featuring this goal.
Model SR SPL
Refer to caption DEBiT-b 40.5 17.1
Refer to caption RANa-b 46.5 17.0
Figure 5: Navigation in simulation from real goal images, retrieving new goals: 2 episodes shown, 200 episodes evaluated.
5.4 Further ablation studies
All ablations were conducted on RANa-t models based on the DEBiT-t backbone to allow for faster training and evaluation. Additional ablations are reported in Supplementary.
Model Retrieval Encoder Retr. Feature SR SPL
[Uncaptioned image] DEBiT-t 79.9 51.4
[Uncaptioned image] RANa-t DINOv2 Gumbel CroCo Geom. FM 83.0 52.3
[Uncaptioned image] RANa-t DINOv2 Gumbel hwRN18 79.1 48.2
Table 7: [Uncaptioned image]  Impact of Geometric FM on context representation.
Geometric FM provides essential directional information – and plays a crucial role in the effectiveness of RANa. We validate this intuition by training with the same procedure as all other models a RANa-t ImageNav network where the encoders $d$ in Eq. 6 are replaced by a hwRN18 (same as encoder $x$) that takes as $9{\times}112{\times}112$ input the channel-concatenated observation, goal and context, and outputs the context features $\mathbf{e}_{t,n}$, fed to the Gumbel selector. As shown in Tab. 7, this model achieves $79.1$ SR, against $83.0$ SR of RANa-t: in this setting, the agent is not capable of exploiting context information and performance remain at DEBiT-t baseline level, validating the importance of the geometric pre-training done with the CroCo FM in [6].
Retrieval is efficientTab. 8 shows model size, retrieval ($t_{R}$) and inference ($t_{I}$) time for selected DEBiT and RANa models. Retrieval time is low (negligible in non-graph case), could even be further optimized by not retrieving at every step, and scales favorably with the dataset size thanks to efficient approximate nearest neighbor search libraries such as Faiss [22]. Inference time is slightly higher for RANa compared to DEBiT because of encoder $c$ that computes CroCo features for context images, observation and goal. Still, RANa-t is smaller and faster than DEBiT-b while achieving similar performance, suggesting that retrieval augmentation can be a frugal alternative to model scaling.
Context diversity is importantTab. 9 evaluates the behavior of RANa-t ImageNav model when using different strategies to create the context, (i) rand: 8 random elements from the retrieval dataset, (ii) top 8: most similar database images to the goal in DINOv2 feature space, and (iii) top 8 MMR: most similar database images filtered by MMR as in Sec. 4.1. Filtering the context with MMR is important, as the agent relies on diverse context items; removing MMR degrades performance. Other approaches have lower but more constant performance across settings. Similar experiments and behavior is presented for ObjectNav in Supplementary.
Model Retrieval Nr Par $t_{R}(s)$ $t_{I}(s)$ SR SPL
DEBiT-t 27M 0.016 79.3 50.0
DEBiT-b 66M 0.025 83.0 55.6
RANa-t DINO-Gumbel 32M 0.001 0.021 83.0 52.3
RANa-t DINOGraph-Attn 34M 0.055 0.024 84.7 63.7
Table 8: Retrieval and inference runtimes (Tesla-T4-16G GPU).
rand top 8 top 8 MMR
rand 81.9 (51.3) 80.3 (50.4) 81.0 (50.3)
top 8 78.4 (49.0) 81.4 (48.1) 80.7 (51.1)
top 8 MMR 77.0 (47.1) 73.7 (42.4) 83.0 (52.3)
Table 9: [Uncaptioned image]  Impact of retrieval data selection - SR (SPL) for ImageNav with different context building strategies, combinations of train and test conditions.
6 Conclusion
We propose a general retrieval-augmented navigation agent, trained with RL, that is able to retrieve and act on images stored in a large, global database of robot observations. We leverage pre-trained foundation models to allow the agent to query and process contextual information that can be useful for solving its target task. Augmenting ImageNav and ObjectNav agents with a context retrieval module improves navigation performance in both tasks, despite the goal specifications by different modalities (an image in ImageNav , text in ObjectNav), which showcases the generality of the proposed method. Moreover, we show how to use retrieval across domains and modalities to apply existing agents to new tasks in a zero-shot fashion.
This work opens the possibility of letting navigation agents benefit from the large amount of unstructured data collected during the operation of robot fleets. Future work will explore the integration of metadata to the database and data-driven query mechanisms that allow an agent to target the information it needs.
Appendix
Appendix A Implementation details
A.1 Retrieval dataset creation
Both ImageNav and ObjectNav benchmarks use the Habitat simulator, platform and task definitions [64]. For ImageNav, we use the Gibson dataset [77], consisting of 72 train and 14 eval scenes, and the ImageNav agent configuration: height 1.5m, radius 0.1m, $112{\times}112$ RGB camera with $90^{\circ}$ horizontal field-of-view (hfov) at 1.25m from the ground. For ObjectNav, we use 80 train and 20 eval scenes of the HM3DSem-v0.2 [60, 81] as in [10] and an agent of height 0.88m, radius 0.18m with $112{\times}112$ RGB sensor with $79^{\circ}$ hfov placed at 0.88m from the ground. For each scene we generate a retrieval dataset $\mathcal{D}$ usually containing $1,000$ FPV images as follows: (i) we randomly choose start and goal viewpoints such that the goal is navigable from the start, (ii) we make an agent follow the shortest path from start to goal, and (iii) after each action, we save the agent’s RGB observation as an image in $\mathcal{D}$. We repeat this process until the dataset contains the desired number of images. We randomly choose the distribution of starts and goals such that the agent roughly covers all navigable areas, and that the path between pairs of viewpoints are not too close to one another to simulate a prior offline exploration of the scene.
A.2 RANa agent architecture
In this section, we further explain some of the architecture choices made for our method. Observation encoderx processes the RGB input image $\mathbf{x}_{t}\in\mathbb{R}^{3\times 112\times 112}$. It is implemented with a half-width ResNet-18 [31] and generates an output feature $\tilde{\mathbf{x}}_{t}\in\mathbb{R}^{512}$.
Goal comparison functiong compares the observation $\mathbf{x}_{t}$ to the goal $\mathbf{g}$. For ImageNav we use the frozen binocular encoder of DEBiT [6], based on the Geometric Foundation Model (FM) CroCo [74]. This encoder is a Siamese encoder $E$ applied to both $\mathbf{x}_{t}$ and $\mathbf{g}$, and a decoder $D$ which combines the output of the two encoders. Both $E$ and $D$ are implemented with a ViT architecture with self-attention layers, and $D$ also adds cross-attention. The output of the decoder $D$ is further compressed by a fully connected layer (frozen, from the DEBiT model) that projects the flattened output of $D$ into $\tilde{\mathbf{g}}_{t}\in\mathbb{R}^{3136}$. For details about the binocular encoder $g$ we refer to [6] and [74].
In ObjectNav we use OpenCLIP [35] to encode the visual observation $\mathbf{x}_{t}$ and the textual goal $\mathbf{g}$. This process produces two features $cl_{V}(\mathbf{x}_{t})\in\mathbb{R}^{512}$ and $cl_{T}(\mathbf{g})\in\mathbb{R}^{512}$, which are concatenated to generate the output $\tilde{\mathbf{g}}_{t}\in\mathbb{R}^{1024}$.
Action embeddingl encodes the previous action $\mathbf{a}_{t-1}$ into a $32$-dimensional feature vector $\tilde{\mathbf{a}}_{t}$ and it is taken from the DEBiT [6] model. For ImageNav, we load DEBiT action encoder and freeze it during training, while for ObjectNav, we trained it from scratch, since the two tasks do not share a common action space. Indeed, the task definition of ObjectNav requires two additional actions, LOOK UP and LOOK DOWN.
Gumbel context encoderVariant 1 of the trainable context encoder c (Sec. 3.1 ) selects one element $\tilde{\mathbf{c}}_{t}$ among the context features $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ based on a learned relevance estimate $\{\alpha_{n}\}_{n=1}^{N}$. The retrieved context images are compared to the goal image (if available) and to the observation by the same frozen binocular encoder of DEBiT described above, leading to the $N$ context features $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ each of dimension $\mathbb{R}^{D}$, where $D=6272$ for ImageNav or Instance-ImageNav (when goal image is available), and $D=3136$ for ObjectNav. A fully connected layer $\texttt{Linear}(D,1)$ (using a syntax inspired by PyTorch) estimates the relevance $\{\alpha_{n}\}_{n=1}^{N}$ for each context element and the item with largest relevance is selected using the Gumbel soft-max sampler [37]. In the ImageNav or Instance-ImageNav cases, only the binocular feature representing the comparison between observation and context item is selected, so in all cases the context feature $\tilde{\mathbf{c}}_{t}\in\mathbb{R}^{3136}$.
Attention-based context encoderVariant 2 of the trainable context encoder creates a context feature by computing the self-attention over the context elements $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ computed as described in the previous paragraph. In this case the context encoder c first reduces the dimensionality of each context feature to $\mathbb{R}^{128}$ through a fully connected layer $\texttt{Linear}(D,128)$ ($D=6272$ for ImageNav or Instance-ImageNav, and $D=3136$ for ObjectNav), then processes them with a standard two-layer transformer encoder network [72] with $4$ attention heads and feedforward dimension $256$. The $N$ context features are concatenated into a feature vector of size $\mathbb{R}^{N\times 128}$ and processed by a 2-layer $\texttt{MLP}(128,3136)$ with hidden dimension $128$ and output $3136$, such that the context encoder output $\tilde{\mathbf{c}}_{t}$ has the same dimension of that of the Gumbel context encoder.
Recurrent memory – The flatten vectors $\tilde{\mathbf{x}}_{t}$, $\tilde{\mathbf{g}}_{t}$, $\tilde{\mathbf{a}}_{t}$ and $\tilde{\mathbf{c}}_{t}$ are concatenated and fed to a single-layer GRU [18] with hidden state $\mathbf{h}_{t}\in\mathbb{R}^{512}$. This GRU is an “extended” version of the baseline agent’s GRU without context, which is finetuned. However, since RANa concatenates the context feature $\tilde{\mathbf{c}}_{t}$ to the other inputs, we cannot directly finetune the GRU as is. We modify the GRU of the base model by padding the input weight matrices of the first layer, $W^{0}_{ir},W^{0}_{iz},W^{0}_{in}$ (and corresponding bias vectors $b^{0}_{ir},b^{0}_{iz},b^{0}_{in}$) along the input dimension with random values. The remaining GRU matrices are initialized to the values of the baseline checkpoints and finetuned from there.
Policy – The hidden state $\mathbf{h}_{t}$ is fed to a linear policy $\pi$ composed of a linear Actor head that generates a softmax distribution over actions and a linear Critic head that evaluates the current state.
A.3 DINOv2-Graph
In this section we provide details about the construction of the DINOv2-Graph context. The graph is built from the affinity matrix $\Omega_{\mathcal{D}}$ containing the similarity between DINOv2 [55] features of all pairs of images in $\mathcal{D}$, where the nodes are images $\mathbf{x}_{i}^{\mathcal{D}}\in\mathcal{D}$ and edges are weighted by similarity. We use Dijkstra’s algorithm [20] (from the SciPy python package) to select the shortest path between two nodes.
Figure 6: Examples of shortest paths on DINOv2-Graph (SWG) and on Pose Graph (PS) - The most similar image $r_{1}^{\mathbf{x}_{t}}\in\mathcal{D}$ to the goal is highlighted in red and localized with a red star on the floorplan, while $r_{1}^{\mathbf{g}}\in\mathcal{D}$, in blue, is the most similar image to the current observation. Green crosses connects images along the shortest-path in the DINOv2-Graph (corresponding images are highlighted in green on top of the map) and purple crosses refer to poses on the Pose Graph, with corresponding images shown with purple contour at the bottom of the map.
Selecting a path on this semantic graph (DINOv2-Graph) is not necessarily equivalent to a “physical path” (see examples in Fig. 6). In order to enforce selection of edges between images that have high overlap probability (i.e.represent the same part of the scene) we made two further modifications to the above graph. First, as we observed that a DINOv2 similarity score above $th=0.75$ between L2 normalized features reflects, in most cases, a good overlap between the two images, we only kept edges for which the similarity was above this threshold, removing all the others. Second, in order to enforce Dijkstra’s algorithm to select paths with high similarity (low cost), we used the weights $w_{ij}=\sqrt{(1-s_{ij})}$, where $s_{ij}$ is the DINOv2 similarity between the L2 normalized features.
During both training and inference, at each step, the agent finds in $\mathcal{D}$ the most similar images $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ to the observation $\mathbf{x}_{t}$ and goal $\mathbf{g}$, respectively, and computes the shortest path between $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ on the DINOv2-Graph. Let $\mathcal{P}=\{r_{1}^{\mathbf{x}_{t}},r_{1},r_{2},...r_{n},r_{1}^{\mathbf{g}}\}$ be the set of nodes on the shortest path and $C$ the context size. If $n+2>C$, we sample images randomly from $\mathcal{P}$ and update all images in the context. If $n+2\leq C$, we use all images in $\mathcal{P}$ to update $n+2$ images in the context. If no path was found, $n=2$ hence we only update two elements of the context with $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$.
SWG SBG DWG PG
RANa-b (Gumbel) 89.7 (60.4) 89.6 (60.5) 88.7 (60.7) 89.7 (60.7)
RANa-b (Attention) 89.8 (71.0) 89.6 (70.3) 90.7 (71.8) 91.9 (69.8)
RANa-t (Attention) 84.7 (63.7) 83.5 (61.1) 86.8 (64.8) 87.4 (62.2)
Table 10: [Uncaptioned image]  Impact of the graph structure on RANa with DINOv2-Graph - SR (SPL) for ImageNav with retrieval using different graph structures at test time. All RANa models were trained with Sparse Weighted Graph (SWG). In bolded best results the PG variant is not considered as it uses the camera poses of images in $\mathcal{D}$, we only add it for comparison.
Analyzing the choice of the graph structure – We trained our models on the Gibson dataset where we had around $10$k images per scene in the retrieval dataset $\mathcal{D}$. We selected the checkpoint obtained at $130$M steps for all models. During testing we used a set of $|\mathcal{D}|=1$k images per scene. In this section we study how the graph structure affects performance. The default model, used for the results reported in Tab. 2 of the main paper, builds a weighted graph by removing edges with similarity below $th=0.75$ and uses $w_{ij}=\sqrt{(1-s_{ij})}$ as weights. We call this Sparse Weighted Graph (SWG). We evaluate three additional strategies to construct the DINOv2-Graph. (i) We consider the same graph, but instead of using $w_{ij}=\sqrt{(1-s_{ij})}$ we use $w_{ij}=1$ as edge weights, and we call this model Sparse Binary Graph (SBG). (II) We also consider a similarity threshold $th=0.4$ which yields a much denser graph, called Dense Weighted Graph (DWG). (iii) Finally, we compare these results with a “pose graph”, where instead of using DINOv2 features, we use distances between camera poses to build the affinity map. In this case we consider a binary graph where we cut edges for which the distance between the camera poses is above $1$m. We denote the Pose Graph with PG. From the results shown in Tab. 10 we can draw the following observations. First, using a pose aware graph (PG) does not necessarily improve the accuracy over using DINOv2-Graph, and when it does, the improvement is rather small. Note however that the pose of the observation and goal images were not used to select the start and end points of the shortest path, $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$, these were retrieved using DINOv2 feature similarity and we only replaced the DINOv2-Graph with the Pose Graph. The second observation we can make is that the graph structure does not have a big impact on the model with a Gumbel selector, but a bigger one on the model with attention-based context encoder. This is not surprising as the Gumbel selects an image from the context and use it independently from other context elements, while Attention combines information from all images within the context, hence it is more important how we fill up the context. Finally, we observe that the Dense Weighted Graph (DWG) performs better than the Sparse Weighted Graph (SWG). We conjecture that the reason is that the model was trained with $\mathcal{D}$ of size $10$k images and we rarely had no connection between two requested nodes ($r_{1}^{\mathbf{x}_{t}}$ , $r_{1}^{\mathbf{g}}$), which occurs more often at test when only $1$k images are available. Using a denser graph (DWG) probably helps overcoming these cases.
Similarities and differences with topological maps – The proposed graph representation is related to topological map structures [63, 4, 14, 76, 40, 69]. However there are several critical differences between our model and traditional topological map based methods. First, we do not use the topological map to localize the agent, plan, and act to reach the “next waypoint”. Instead, the information gathered from planning is passed to the agent as a “recommendation” in the form of the retrieval context, to be integrated into decision making through learning: our agent is a full fledged policy which has access to memory, goal and a series of images along the path, which arguably enables more informed decision-making, and improved robustness and navigation performance. Additionally, our graph is connected based only on visual similarity rather than predictions from temporal distances. While it is debatable whether this approach is inherently superior, it offers a distinct advantage: we do not require trajectories to build the graph but can build it from a large number of isolated images. The main drawback is that we rely on the assumption that visual proximity is a proxy for spatial proximity, which is not always true (see Fig. 6). Future work could attempt to learn to take into account the uncertainty in this assumption through data-driven planning, as in [4].
Appendix B Additional ablation experiments
In this section, we provide additional ablation studies that could not fit into the main document. As for the experiments in Sec. 5.4 of the main paper, all experiments are conducted on RANa-t models with Gumbel selector as context encoder $c$.
4 8 12 16 20 24
4 80.8 (52.3) 80.5 (51.6) 80.4 (52.7) 81.0 (52.3) 80.9 (52.0) 81.7 (54.0)
8 80.4 (49.8) 83.0 (52.3) 80.6 (50.6) 84.0 (53.0) 82.0 (53.2) 82.9 (53.4)
12 81.9 (50.0) 80.0 (49.2) 81.9 (51.2) 80.6 (50.6) 81.6 (50.8) 82.4 (52.2)
Table 11: [Uncaptioned image]  Impact of context size for RANa-t - SR (SPL) for ImageNav with retrieval over different context sizes, combinations of train and test conditions.
4 8 12 16 20 24
4 51.6 (26.1) 50.8 (25.7) 49.8 (25.3) 49.8 (25.3) 49.0 (24.7) 50.6 (25.8)
8 50.5 (25.9) 51.9 (26.4) 52.1 (26.9) 50.3 (26.6) 51.6 (26.5) 51.1 (26.6)
12 51.4 (26.4) 51.4 (26.9) 50.9 (26.7) 51.2 (27.3) 50.9 (26.7) 51.6 (27.4)
Table 12: [Uncaptioned image]  Impact of context size for RANa-t - SR (SPL) for ObjectNav with retrieval over different context sizes, combinations of train and test conditions.
Impact of context size – is studied in Tab. 11 for ImageNav ([Uncaptioned image]) and in Tab. 12 for ObjectNav ([Uncaptioned image]). One interesting advantage of the Gumbel selector is that it allows for context of different sizes during evaluation, independent of the context size used during training. With that said, this parameter has a relatively minor impact on performance, with larger training context sizes showing positive performance trends. The size $N=8$ is a choice that works well both for ImageNav and ObjectNav and balances navigation performance with execution time. For ObjectNav, we observe a more pronounced performance increase for larger context sizes during evaluation, so we use $N=8$ during training and $N=12$ during evaluation.
[…]
Model Context SR SPL
RANa-t top 8 83.0 52.3
RANa-t top 1 79.6 48.7
Figure 7: Refer to caption  Histogram of the rank of the context element selected by RANa-t with Gumbel context encoder on the ImageNav Gibson eval set.
Impact of retrieval database size – on zero-shot Instance-ImageNav ([Uncaptioned image]) is shown on Tab. 13. Here the goal image $\mathbf{g}^{\mathcal{R}}$ is retrieved from the database, and thus retrieval accuracy is crucial for success and avoid terminating an episode too far from the actual goal. Tab. 13 shows this phenomenon, with zero-shot Instance-ImageNav performance clearly increasing with retrieval dataset size.
1k 5k 10k 20k 50k 100k
47.6 (20.2) 56.2 (24.4) 56.4 (23.8) 57.4 (25.0) 58.7 (26.8) 59.1 (25.3)
Table 13: [Uncaptioned image] Impact of retrieval database size in zero-shot - SR (SPL) in Instance-ImageNavigation increases with dataset size.
Analysis of “oracle” context – for ImageNav ([Uncaptioned image]) is carried out in Tab. 14. In this experiment we study which information is most useful and can be exploited by RANa, by building contexts using privileged information in simulation – therefore in these cases the retrieval dataset $\mathcal{D}$ is not used. Oracle 1 captures a panoramic view of the goal location and is created by collecting $N=8$ context images at the goal position while rotating the camera view by $45^{\circ}$ clock-wise for each element. This context represents a soft upper bound for the static context in Sec. 4.1 , where a fixed set of items is chosen per episode based on their similarity with the goal image. Oracle 2 context is richer and it is generated by computing the shortest path from the agent position to the goal at each step and collecting views along this path. This is a soft upper bound for the DINOv2-Graph context (Sec. 4.2 ), which is dynamically generated at each step on the context graph. Tab. 14 shows that, as expected, both types of oracle context can provide large gains, especially in SPL, i.e.navigation efficiency. The Table suggests that there is little room for improvement in success (SR) using static goal context, while more structured information, e.g.paths, derived from data in $\mathcal{D}$ can improve navigation considerably.
Context Type SR SPL
RANa-t goal static 83.0 52.3
Panorama (oracle 1) static 83.4 71.6
RANa-t DINOv2-graph dynamic 84.7 63.7
Shortest path (oracle 2) dynamic 98.3 87.9
Table 14: [Uncaptioned image]  Navigation with “oracle” context - SR (SPL) for ImageNav using privileged information to build oracle context.
Impact of using pre-trained GRU weights – As mentioned earlier all RANa agents use a GRU finetuned from the baseline agent’s GRU. We evaluate the impact of this choice for the ImageNav ([Uncaptioned image]) model by training the RANa-t GRU from scratch while keeping all other settings fixed. Tab. 15 shows that finetuning the GRU form pretrained weights – those of DEBiT [6] in this case – improves navigation performance. We also observe faster convergence during training, which might in part explain the performance increase.
Model SR SPL
GRU finetuned 83.0 52.3
GRU from scratch 81.4 51.0
Table 15: [Uncaptioned image]  Impact of GRU finetuning - SR (SPL) for ImageNav with GRU finetuned and trained from scratch.
Context selection criteria – While in RANa ObjectNav the context selection is always based on the observation, as the goal image is not available, in RANa ImageNav model we have both observation and goal images. All ImageNav experiments reported in the paper use a feature vector for each context element which is obtained by concatenating two features: (i) the output of the Geometric FM $d$ comparing context element and observation, and (ii) the output of $d$ comparing context element and goal. We evaluate the impact of this choice by training a RANa-t model which selects context elements based only on feature (i) which compares context elements to the observation. The performance of this variant, compared to the baseline which uses both observation and goal, is displayed in Tab. 16, which shows that information provided by comparing the context to the goal is useful and can be exploited by RANa to navigate more successfully and efficiently to the goal.
Model SR SPL
context selection based on obs and goal 83.0 52.3
context selection based on obs only 81.0 50.3
Table 16: [Uncaptioned image]  Impact of context selection - SR (SPL) for ImageNav with context elements selected based on comparisons with only the observation, or both goal and observation.
Context building for ObjectNavTab. 17 shows the behavior of RANa-t ObjectNav ([Uncaptioned image]) when adopting different context building approaches, (i) rand: 8 random elements from the retrieval dataset, (ii) top 8 $cl_{S}$: most similar database images to the object goal encoder in CLIP feature space using CLIP similarity as metric, and (iii) top 8 $\omega_{\mathbf{o}}$: most similar database images using the score softmax normalization described in Sec. (4.1 ) of the main paper. The observed behavior is similar to ImageNav (Sec. 5.4 ). First, a RANa-t model trained with a context with randomly selected images from the scene gallery performs worse than a model trained with either top 8 images. Nevertheless, this model still outperforms our baseline trained without retrieval (SR=$42.9$ and SPL=$23.0$ for the non-retrieval baseline) and is more robust to random context images during evaluation. For RANa-t models trained with retrieval top 8 context, we observe the improved performance of using softmax normalized scores.
rand top 8 $cl_{S}$ top 8 $\omega_{\mathbf{o}}$
rand 46.8 (23.8) 47.7 (25.0) 46.1 (23.1)
top 8 $cl_{S}$ 44.3 (23.0) 48.0 (25.4) 48.5 (26.0)
top 8 $\omega_{\mathbf{o}}$ 44.4 (21.4) 51.3 (25.9) 51.9 (26.4)
Table 17: [Uncaptioned image]  Impact of retrieval data selection - SR (SPL) for ObjectNav with different context building strategies, combinations of train and test conditions.
Impact of retrieval dataset size for ObjectNav – is shown in Tab. 18. Results demonstrate that unlike ImageNav (Tab. 5 ), ObjectNav seems to need larger retrieval datasets both at training and at testing. Performance degrades significantly when the dataset $\mathcal{D}$ at training only has $1,000$ images. At test, using larger retrieval datasets leads to better performance, with metrics degrading significantly with $\mathcal{D}$ containing $100$ items.
100 1k 10k 50k
1k 43.8 (21.7) 45.3 (22.4) 46.8 (23.2) 47.1 (23.2)
10k 46.6 (24.9) 50.3 (26.4) 52.1 (26.9) 51.8 (27.2)
Table 18: [Uncaptioned image]  Impact of retrieval database size - SR (SPL) for ObjectNav with retrieval using different database sizes, combinations of train and test conditions.
Impact of CLIP variants for ObjectNav – is evaluated in Tab. 19. In this experiment, we build retrieval datasets $\mathcal{D}$ of size $10$k images for each of the 100 HM3DSem-v0.2 scenes used for ObjectNav (80 train, 20 test). For each of the 6 target object categories (chair, bed, plant, toilet, tv monitor, sofa) and for each scene, we do a search in the retrieval dataset $\mathcal{D}$ by querying the CLIP-encoded goal category (i.e.the CLIP text encoder for the sentence “a picture of a category inside a house”) and comparing it with CLIP encoded dataset images of the scene, providing top 1, top 8 and top 20 success percentages. A retrieval is considered successful if the retrieved image (whose ground-truth pose is known), is closer to one of the viewpoints of a object instance of the correct category than the success threshold of the ObjectNav task, i.e.1111m. The Table shows that OpenCLIP performs better than alternative CLIP models, namely OpenAI’s original CLIP weights [59] and SigLIP [85], all with the softmax normalization proposed in Sec. 4.1 being used. As observed earlier, ranking images by CLIP similarity to the category does not perform well [58]; since the number of target classes is known (6 in all our experiments), rescaling the score of each image in $\mathcal{D}$ via softmax significantly boosts performance.
Vision-language model softmax (Sec. 4.1 ) top 1 top 8 top 20
OpenCLIP [35] 40.6 76.0 88.5
OpenCLIP [35] 78.1 95.8 97.9
CLIP [59] 67.7 88.5 90.6
SigLIP [85] 70.8 91.7 94.8
Table 19: Retrieval performance of CLIP variants in the context of ObjectNav ([Uncaptioned image]). Retrieval is considered success if the retrieved image is of the correct category and if it is closer than $1$m to the target object.
Appendix C Qualitative analysis
C.1 Examples of retrieved context
In this section we provide visual examples of the different context types used by our RANa model.
Fig. 8 shows three types of context used in this work for RANa ImageNav ([Uncaptioned image]), for $N=8$: top 8 most similar dataset images to the image goal in DINOv2 feature space (top) and top 8 MMR most similar images filtered by Maximum Margin Relevance (MMR) (middle), are static context types, i.e.they are fixed along one episode and depend only on the goal image, shown on the right. DINOv2-Graph (bottom) is a dynamic context that is recalculated at each step and depends on the observation (left) as well.
This example clearly shows that the most similar images to the goal are very similar to each other as well, and might not be very useful to the RANa agent. Filtering images using MMR increases the diversity of the context items, and this, as shown in Tab. 9 in the main text, increases RANa performance. For navigation purposes, the DINOv2-Graph context seems the most informative, as it provides waypoints from the current location to the goal – indeed this context provides the best navigation performance for RANa, see Tab. 2 .
Figure 8: Refer to caption  Context types for ImageNav - From top to bottom: top 8 most similar dataset images to the image goal (right), top 8 MMR most similar images filtered by MMR, and DINOv2-Graph, the only one depending on the observation (left) as well.
Fig. 9 displays examples of context for RANa ObjectNav ([Uncaptioned image]) using similarity in CLIP feature space for the six different goal categories, chair, bed, plant,sofa, tv monitor and toilet. As discussed in the paper, results accurately represent instances of the target object present in the scene, with the exception of the tv monitor class where we observe some errors, especially in scenes where monitors are absent.
Figure 9: Refer to caption  Examples of context for different ObjectNav categories.
C.2 Navigation examples
Figure 10: Refer to caption  RANa-b successfully navigates to challenging goals in two ImageNav episodes - The agent observation is on the left, next to it is the goal image highlighted in green, then there are $N=8$ context elements, with the one selected by the RANa agent at the current time step highlighted in red, and on the right the agent path drawn on the floorplan of the scene, with the goal position indicated by a red square. At the bottom we show the trajectory of the baseline DEBiT-b agent on the same episode.
In Fig. 10 we show examples of our agent RANa-b navigating to two challenging ImageNav goals. The figures visualize the observation (left), goal (highlighted in green), $N=8$ context elements, with the one selected at the given time step highlighted in red, and the agent path drawn on the scene floorplan on the right. Each row of pictures focuses on one time step along the episode and are ordered in time from top to bottom. In this figure and the following ones, time is not sampled uniformly: we selected time steps, typically towards the end of the episode, where the agent makes interesting context selections to solve the episode.
The episode in Fig. 10 (left) is challenging because the environment is rather complex, the goal image is ambiguous and it is “hidden” in a terrace at the extreme perifery of the navigable space. The baseline DEBiT-b agent, without context, does its best to explore the apartment and gets close to the terrace door where the goal is located, but does not manage to navigate to it. RANa-b selects from the context an image (highlighted in red) that contains the railing of the terrace that helps our agent solve this complex navigation episode. The navigation episode on the right is complex for similar reasons: large complex environment and goal image with very little information. RANa is capable of exploiting the few relevant context items, by first selecting the context image containing the door to the room where the goal is located, and then the cabinet, whose corner is visible on the goal image.
Figure 11: Refer to caption  RANa-b successfully navigating to two goals in zero-shot Instance-ImageNav episodes - The agent observation is on the left, next to it is the original goal image from the Instance-ImageNav task (in cyan) and the retrieved goal image $\mathbf{g}^{\mathcal{R}}$ highlighted in green. The remaining data is organized and visualized as in Fig. 10.
Fig. 11 shows RANa-b ImageNav agent successfully navigating in two zero-shot Instance-ImageNav episodes. Results are visualized in the same way as in Fig. 10, with the exception of the goal image. In this case we have in fact two goal images, one from the original task (highlighted in cyan) which is taken with a sensor with different pose and intrinsics of the agent sensor, and the goal image $\mathbf{g}^{\mathcal{R}}$, highlighted in green, which is retrieved from the scene database $\mathcal{R}$. In both cases, RANa selects context items that represent elements visible in our agent observation (the toilet on the left and the door on the right) that guide the agent inside the room where the goal is located. Interestingly, in both cases the baseline agent passed along the same path, but did not enter the room, arguably because it did not have access to the context.
Figure 12: Refer to caption  RANa-b successfully navigates to real-world goals in two zero-shot Instance-ImageNav episodes - In this case the original image goal (highlighted in cyan) is taken with a smartphone in a real environment. Again we retrieve a goal image $\mathbf{g}^{\mathcal{R}}$ from $\mathcal{D}$ and the agent navigates in a 3D model of the environment.
In Fig. 12 we present an even more challenging scenario, where RANa navigates in two zero-shot Instance-ImageNav episodes, where the goal image is captured with a smartphone in a real environment. As mentioned in the main text, we have access to this environment and a 3D model of the scene is available. The agent navigates to the retrieved goal image $\mathbf{g}^{\mathcal{R}}$ from the database $\mathcal{D}$ in the simulated reconstruction of the scene. The example on the left is particularly interesting because the retrieved goal image, in green, is not the correct instance of the original goal. However, guided by several context items that contain instances of the original goal image, RANa succeeds in reaching the goal. In the navigation episode on the right, our agent picks relevant context elements containing the chairs of the meeting room where the goal is positioned, and enters the room solving the navigation episode.
Figure 13: Refer to caption  A failure case for RANa-b - Sometimes context degrades the performance of the baseline agent. This example shows how RANa reaches the zone of the goal image and then starts turning around, circling through context items. In contrast DEBiT-b (bottom panel) quickly navigates to the goal.
Finally, we observe cases in which access to a context deteriorates the performance of the baseline agent. We can describe the behavior of our RANa agent in these cases as being “distracted” by context items visible by the agent at a given time step. Fig. 13 shows a typical instance of this phenomenon: the baseline DEBiT agent (trajectory shown at the bottom) quickly gets to the goal, while RANa reaches the goal area, but starts turning on itself, cycling multiple times through relevant, but “distracting”, context images.

### v2

@@
-Gianluca Monaci          Rafael S. Rezende          Romain Deffayet          Gabriela Csurka Guillaume Bono          Hervé Déjean          Stéphane Clinchant          Christian Wolf
+RANa: Retrieval-Augmented Navigation
 Abstract
@@
 We address this by introducing a new retrieval-augmented agent, trained with RL, capable of querying a database collected from previous episodes in the same environment and learning how to integrate this additional context information.
-We introduce a unique agent architecture for the general navigation task, evaluated on ObjectNav, ImageNav and Instance-ImageNav.
-Our retrieval and context encoding methods are data-driven and heavily employ vision foundation models (FM) for both semantic and geometric understanding.
+We introduce a unique agent architecture for the general navigation task, evaluated on ImageNav, Instance-ImageNav and ObjectNav.
+Our retrieval and context encoding methods are data-driven and employ vision foundation models (FM) for both semantic and geometric understanding.
 We propose new benchmarks for these settings and we show that retrieval allows zero-shot transfer across tasks and environments while significantly improving performance.
+Refer to caption(a) agent(b) use cases(c) inference runtime
+Figure 1: Retrieval-Augmented Navigation (RANa).
+(a) We augment a navigation agent with context information retrieved from a scene-specific retrieval database, fed with data by one or potentially multiple robots.
+We leverage semantic or multi-modal foundation models for retrieval, and geometric foundation models for context encoding.
+(b) We tackle ImageNav, Instance-ImageNav and ObjectNav tasks, demonstrating performance improvements (Success +x), opening new use cases from existing models, and zero-shot applications with retrieval only added at test time (Refer to caption).
+(c) Retrieval is efficient, neglectable compared to vision models and keeps inference run-time in the same ballpark as non-augmented models like DEBiT.
 1 Introduction
@@
 Most modern methods tackling this scenario, e.g.
-[75, 12, 79, 6], use learning-based approaches in an episodic setting: every episode in the robot’s life is treated as if it was its first.
+Wijmans et al.
+(2019); Chaplot et al.
+(2020a); Yadav et al.
+(2023a); Bono et al.
+(2024a), use learning-based approaches in an episodic setting: every episode in the robot’s life is treated as if it was its first.
 This is obviously sub-optimal: during operation the scene is explored, objects and their locations are observed, affordances found, failure cases encountered etc.
 Exploiting this information is crucial for designing efficient navigation agents that can continuously improve.
-Figure 1: We augment a navigation agent with context information retrieved from a scene-specific retrieval database, fed with data by one or potentially multiple robots.
-We leverage semantic and multi-modal foundation models for retrieval, and geometric foundation models for context encoding, harnessing its power to provide directional information.
-We explore ObjectNav, ImageNav and Instance-ImageNav tasks, including zero-shot cases.
-Benchmarks like the k-item scenario [4], multi-object navigation [73] or GOAT benchmark [39] test the capacity of an agent to retain information from previous (sub-)goals and trajectories.
-Structured latent memory trained end-to-end with RL performs best in these benchmarks [50].
-However, these architectures have only been tested on short (multi-)episodes of $500$ to $2,500$ agent steps and are hardly suitable for a real continuous operation: transformer-based agents with self-attention over time suffer from limited context length and the quadratic complexity of attention, while recurrent models are limited by the size of their latent memory and the network capacity growing quadratically with representation size.
-Alternative ways to integrate information from initial rollouts are topological maps [63, 69], which require the estimation and generation of a structured graph model from a sequence of posed observations.
+Benchmarks like the k-item scenario (Beeching et al., 2020b), multi-object navigation (Wani et al., 2020) or GOAT benchmark (Khanna et al., 2024) test the capacity of an agent to retain information from previous (sub-)goals and trajectories.
+Structured latent memory trained end-to-end with RL performs best in these benchmarks (Marza et al., 2022).
+However, these architectures have only been tested on short (multi-)episodes of 500500500 to 2,5002,5002 , 500 agent steps and are hardly suitable for a real continuous operation: transformer-based agents with self-attention over time suffer from limited context length and the quadratic complexity of attention, while recurrent models are limited by the size of their latent memory and the network capacity growing quadratically with representation size (Jose et al., 2018).
+Alternative ways to integrate information from initial rollouts are topological maps (Savinov et al., 2018; Sridhar et al., 2024), which require the estimation and generation of a structured graph model from a sequence of posed observations.
 The extension of these methods to multi-robot scenarios is anything but trivial.
-In this work we propose a simple data-driven approach to extend a state-of-the-art navigation agent [6] working out-of-the-box, with capabilities to store visual observations in a global indexed database, share them with a fleet, and retrieve and process them to improve performance during operation.
-The proposed architecture is general and can be used for different navigation tasks — we evaluate our work on ObjectNav, ImageNav and Instance-ImageNav.
+In this work we propose a simple data-driven approach to extend a state-of-the-art navigation agent (Bono et al., 2024a) working out-of-the-box, with capabilities to store visual observations in a global indexed database, share them with a fleet, and retrieve and process them to improve performance during operation.
+The proposed architecture is general and can be used for different navigation tasks — we evaluate our work on ImageNav, Instance-ImageNav and ObjectNav.
 While the goal may be specified either as an image or a semantic category, depending on the task, the retrieved information is always visual first person views (FPVs).
-Computer vision is at the heart of the agent in multiple ways: data-driven retrieval mechanisms leverage vision foundation models (FM) like CLIP and DINOv2 to retrieve candidate views, while a geometric FM [74] integrates the retrieved context and provides directional information to the agent, cf.
-Fig.
-1.
-This approach is scalable, as the dataset can potentially be very large and can be queried by any number of agents with sub-linear time algorithms.
+Data-driven retrieval mechanisms leverage vision foundation models (FM) like DINOv2 (Oquab et al., 2024) to retrieve candidate views, while a geometric FM, CroCo (Weinzaepfel et al., 2022), integrates the retrieved context and provides directional information to the agent, cf.
+Fig.˜1a.
+Retrieval is also fast and keeps inference time in the same ball park of alternative baselines, cf.
+Fig.˜1c.
+This approach is scalable, as the dataset can potentially be very large and can be queried by any number of agents with optimized sub-linear time algorithms.
 Retrieving from an image database is also considerably simpler and more flexible than using involved memory structures such as metric or topological maps.
-The latter require pose to register observations and synchronization to enable multi-agent data collection: unless pose estimates are absolute (which requires scene preparation), fusing new updates from multiple agents requires a jointly maintained map and distributed algorithms.
+While we can use any suitable vision or multi-modal FM to represent, organize and retrieve images from an unstructured database, the latter approaches require pose to register observations and synchronization to enable multi-agent data collection: unless pose estimates are absolute (which requires scene preparation), fusing new updates from multiple agents requires a jointly maintained map and distributed algorithms.
 We propose the following contributions: (i) new navigation tasks, where the environment is augmented with a retrieval database fed by one or multiple robots; (ii) a unified end-to-end architecture for Retrieval-Augmented Navigation, RANa, for multiple tasks; (iii) data-driven retrieval mechanisms that do not require any meta-data beyond raw FPVs; (iv) mechanisms that leverage retrieved information either for additional context, or to replace the goal on the fly, allowing for zero-shot applications, cf.
-Fig.
-2.
-Figure 2: We target several contributions: perf.
-improvement (Success +x), opening new use cases from existing models, and zero-shot applications with retrieval only added at test time (Refer to caption).
-Figure 3: We propose an architecture for Retrieval-Augmented Navigation architecture – RANa, that addresses all use cases of Fig. 2. Dashed arrows are optional depending on the navigation task.
+Fig.˜1.
 2 Related Work
-Visual navigation – The task of navigation has been addressed in robotics research using mapping and planning [71, 46], which requires solutions for mapping and localization [9, 44, 54], planning [41, 66] and low-level control [26, 62].
+Visual navigation – The task of navigation has been addressed in robotics research using mapping and planning (Thrun et al., 2005; Macenski et al., 2020), which requires solutions for mapping and localization (Bresson et al., 2017; Mur-Artal & Tardós, 2017; Labbé & Michaud, 2019), planning (Sethian, 1996; Konolige, 2000) and low-level control (Fox et al., 1997; Rösmann et al., 2015).
 These methods depend on accurate sensor models, filtering, dynamical models and optimization schemes.
-In contrast, end-to-end models learn deep representations such as flat recurrent states [79, 6], occupancy maps [13], semantic maps [12], latent metric maps [56, 4, 32, 50], topological maps [3, 14, 67], self-attention [16, 23, 25, 61] or implicit representations [51].
-Then, they map these representations into actions, using RL [36, 53, 8], imitation learning [21] or by maximizing navigability [7].
+In contrast to them, end-to-end models aim to learn deep representations such as flat recurrent states (Yadav et al., 2023a; Bono et al., 2024a), occupancy maps (Chaplot et al., 2020b), semantic maps (Chaplot et al., 2020a), latent metric maps (Henriques & Vedaldi, 2018; Parisotto & Salakhutdinov, 2018; Beeching et al., 2020b; Marza et al., 2022), topological maps (Savinov et al., 2018; Beeching et al., 2020a; Chaplot et al., 2020c; Sridhar et al., 2024), self-attention (Fang et al., 2019; Du et al., 2021; Chen et al., 2022; Reed et al., 2022; Zeng et al., 2025) or implicit representations (Marza et al., 2023; Kwon et al., 2023).
+Then, they map these representations into actions, using RL (Jaderberg et al., 2017; Mirowski et al., 2017; Bono et al., 2024c), imitation learning (Ding et al., 2019) or by maximizing navigability (Bono et al., 2024b).
 In this work, our goal is to address end-to-end learning of representations and policies with RL.
-Although this problem is nearly solved for PointNav, where the agent must navigate to a relative GPS coordinate [64, 75], it remains a challenge for other goal specifications.
-ObjectNav, where the goal is an object category, requires semantic understanding of the scene, therefore a large body of research builds a semantic map [27, 12] or leverages an LLM to guide navigation based on a scene’s common sense semantics [83, 10, 43, 82].
-Several approaches address this task in a zero-shot fashion, training their agents without ObjectNav rewards and leveraging CLIP encoders for goal representation [70, 47, 27].
-Recently, Zeng et al.
-[84] tackles ObjectNav with a transformer state encoder on simulated environments.
-In ImageNav, an image is provided as goal, requiring a geometric understanding of the scene.
-A common variant is Instance-ImageNav [42], where the goal image depicts a specific object in the environment and is taken with camera parameters different from those of the agent.
-Recently, the end-to-end model DEBiT [6] achieved state-of-the-art performance in ImageNav leveraging the geometric FM CroCo [74] to estimate relative poses between goal and agent’s observation.
-In this work, we build upon DEBiT by both extending it to ObjectNav and augmenting it with retrieval, naturally extending it to the multi-robot setting and enabling zero-shot ObjectNav and Instance-ImageNav.
-To the best of our knowledge no other method performs all these tasks competently, the closest competitor being [47] which tackles ImageNav and zero-shot ObjectNav, although with considerably lower performance (see Sec.
-5).
-Method and Img Obj Not req.: Easiness of Multiple
-representation Nav Nav pose depth repr. update collectors
-GOAT [39] $\sim$ map reg.
-3D Scene graphs [82] $\sim$ graph+map reg.
-Topo map [63, 14, 67] $\sim$ graph upd.
-Semantic maps [12] $\sim$ map reg.
-Sem. Implicit [51] $\nabla$ upd
-Feature SLAM [54] $\sim$ pose graph
-Dense SLAM [24] $\sim$ map reg.
-RANa store
-To support multiple collectors, a method is either pose free, like RANa, or uses distributed algorithms to synchronize, or absolute pose (and not episodic).
-Table 1: RANa compared to SotA in continual navigation. RANa is a flexible, data-driven solution that can easily exploit data from previous episodes, in particular collected by multiple robots.
+Although this problem is nearly solved for PointNav, where the agent must navigate to a relative GPS coordinate (Savva et al., 2019; Wijmans et al., 2019), it remains a challenge for other goal specifications.
+Here we focus on ImageNav, where an image is provided as goal, requiring geometric understanding of the scene.
+A common variant is Instance-ImageNav (Krantz et al., 2023), where the goal image depicts a specific object in the environment and is taken with camera parameters different from those of the agent.
+Recently, the end-to-end model DEBiT (Bono et al., 2024a) achieved state-of-the-art performance in ImageNav leveraging the geometric FM CroCo (Weinzaepfel et al., 2022) to estimate relative poses between goal and agent’s observation.
+In this work, we build upon DEBiT and augment it with retrieval, naturally extending it to multi-robot setting and enabling zero-shot ObjectNav and Instance-ImageNav.
+To the best of our knowledge no other method performs all these tasks competently, the closest competitor being Majumdar et al.
+(2022) which tackles ImageNav and zero-shot ObjectNav, although with considerably lower performance (see Section˜5).
 Retrieval-augmented control – The idea of retrieving information from a dataset of past experiences has been introduced in earlier works.
-Among them, episodic control methods for RL [5, 57, 45, 29, 30, 33] act on successful experiences by re-employing Q-value estimates, enhancing sample efficiency during training.
-In contrast, we exploit retrieved information from past experiences of a robot fleet to enhance navigation performance and to enable zero-shot task generalization.
+Among them, episodic control methods for RL (Blundell et al., 2016; Pritzel et al., 2017; Lin et al., 2018; Hansen et al., 2019; He et al., 2024; Hu et al., 2021) act on successful experiences by re-employing Q-value estimates, enhancing sample efficiency during training.
+In contrast, we exploit retrieved information from past experiences of a robot fleet to enhance navigation performance and enable zero-shot task generalization.
 Goyal et al.
-[28] augment an RL agent with a retrieval process parameterized as a neural network that has access to a dataset of past trajectories.
-Here instead we take inspiration from [34] that, to our knowledge, was the first work to propose a generic non-parameterized retrieval augmentation for an RL agent with previously learned embeddings.
-However, alleviating the need to pre-train embeddings using RL, we use vision foundation models to process stored and retrieved images.
+(2022) augment an RL agent with a retrieval process parameterized as a neural network that has access to a dataset of past trajectories.
+Here instead, we take inspiration from Humphreys et al.
+(2022) that, to our knowledge, was the first work to propose a generic non-parameterized retrieval augmentation for an RL agent with previously learned embeddings.
+Alleviating the need to pre-train embeddings with RL, we use vision foundation models to process stored and retrieved images.
 This enables zero-shot transfer and allows to adapt the approach to various navigation tasks.
-Retrieval-Augmented Planning [38] uses a memory of successful task executions, comprising plans, actions and observation sequences.
+Retrieval-Augmented Planning (Kagaya et al., 2024) uses a memory of successful task executions, comprising plans, actions and observation sequences.
 Past experiences relevant to the task at hand are leveraged to guide planning with an LLM agent based on task constraints and state similarity.
 Xie et al.
-[78] build a hierarchical topological memory to prompt an LLM for reasoning and goal-finding in a navigable environment.
+(2024) build a hierarchical topological memory to prompt an LLM for reasoning and goal-finding in a navigable environment.
 In contrast, we only store images derived from unstructured observations without any associated metadata.
-Navigation supported by prior rollouts – Other navigation methods exploit previous observations to construct a topological map of the explored scene, and use it to guide navigation [63, 14, 67].
+Navigation supported by prior rollouts – Other navigation methods exploit previous observations to construct a topological map of the explored scene, and use it to guide navigation (Savinov et al., 2018; Chaplot et al., 2020c; Shah & Levine, 2022; Sridhar et al., 2024).
 In contrast, our database is unstructured and we can query it to retrieve goal and context images with simple image retrieval.
-In Tab.
-1 we compare RANa with various navigation methods discussed above in the continual navigation scenario.
+Also, here context images are not used as navigation waypoints, but as recommendations, exploited or not by the agent, leading to navigation performance robust to potentially misleading context, as shown in Section˜5.
+In Table˜1 we compare RANa with selected navigation methods discussed above in the continual navigation scenario.
 Most approaches require depth and pose and/or more involved (metric or topological) map updates, which need at least to register observations.
-In contrast, RANa does not require depth or pose, the database update and retrieval is very simple, and naturally supports multiple data collection robots without the need of synchronization.
-While SLAM-based solutions [24, 54] are well established for continual navigation, with RANa we propose a lightweight data-driven solution that can be easily scaled up to take inputs from multiple robots and can tackle ImageNav, Instance-ImageNav and ObjectNav use cases.
+In contrast, RANa does not require depth or pose, the database update and retrieval is straightforward, and naturally supports multiple data collection robots without the need of synchronization.
+While SLAM-based solutions (Engel et al., 2014; Mur-Artal & Tardós, 2017) are well established for continual navigation, with RANa we propose a lightweight data-driven solution that can be easily scaled up to multiple robots and can tackle ImageNav, Instance-ImageNav and ObjectNav use cases.
+Method and Img Obj Not required: Easiness of Multiple
+representation Nav Nav pose depth representation update collectors
+GOAT (Khanna et al., 2024) $\sim$ map reg.
+3D Scene graphs (Yin et al., 2024) $\sim$ graph+map reg.
+Topo map (Savinov et al., 2018; Shah & Levine, 2022) $\sim$ graph update
+Semantic maps (Chaplot et al., 2020a) $\sim$ map reg.
+Sem. Implicit (Marza et al., 2023) $\nabla$ update
+Feature SLAM (Mur-Artal & Tardós, 2017) $\sim$ pose graph
+Dense SLAM (Engel et al., 2014) $\sim$ map reg.
+RANa store
+In order to support multiple collectors, a method should either be pose free, like RANa, or use distributed algorithms to synchronize, or absolute pose (and not episodic).
+Table 1: RANa compared to SOTA in continual navigation. RANa is a flexible, data-driven solution that can easily exploit data from previous episodes, in particular collected by multiple robots.
 3 Retrieval-augmented agent
 We target navigation in 3D environments, where an agent is tasked to navigate from a starting location to a goal and receives at each timestep $t$ an image observation $\mathbf{x}_{t}$.
-Our method is general and the agent can be trained for a diverse range of tasks, but without loss of generality in this work we focus on ImageNav, where the agent receives an additional image goal and ObjectNav, where the agent receives an object goal which is a category in form of text, cf.
-Fig.
-2.
-In both cases we denote the original goal by $\mathbf{g}$.
+Our method is general and the agent can be trained for a diverse range of tasks, but without loss of generality in this presentation we focus on ImageNav, where the agent receives a goal image $\mathbf{g}$.
+An extension to the ObjectNav task is presented in Section˜6.
 The action space is discrete, $\mathcal{A}$ ={MOVE FORWARD 0.25m, TURN LEFT $10^{\circ}$, TURN RIGHT $10^{\circ}$, and STOP}.
 Navigation is considered successful if the STOP action is selected when the agent is within 1m of the goal position in terms of geodesic distance.
-ObjectNav has two additional actions, LOOK UP and LOOK DOWN.
-Our tasks are based on the ImageNav and ObjectNav tasks of the Habitat simulator and platform [64].
-We extended them with a retrieval dataset $\mathcal{D}$ and a retrieval mechanism $\mathcal{R}$, which will be detailed in Sec.
-4.
+Our set-up is based on the ImageNav task of the Habitat simulator and platform (Savva et al., 2019).
+We extend it with a retrieval dataset $\mathcal{D}$ and a retrieval mechanism $\mathcal{R}$, which will be detailed in Section˜4.
 The dataset $\mathcal{D}=\{\mathbf{x}_{i}^{\mathcal{D}}\}$ contains FPVs indexed by $i$ and stemming from previous episodes or from exploratory rollouts in the scene.
-While it is in principle possible to store additional meta-data associated to data collection, like approximate pose, performed actions, reward, success or value (in an MDP sense, when available), we show in this work that the access to FPVs alone provides rich information to the agent (see Sec.
-4).
+While it is in principle possible to store additional meta-data associated to data collection, like approximate pose, performed actions, reward, success or value (in an MDP sense, when available), in Section˜5 we show that access to FPVs alone provides rich, useful information to the agent.
 The agent queries this dataset $\mathcal{D}$ with the retrieval mechanism $\mathcal{R}$, which can be potentially used at each time step $t$.
 $\mathcal{R}$ uses the goal $\mathbf{g}$, and optionally the current observation $\mathbf{x}_{t}$, to retrieve a set of FPVs $\mathcal{R}(\mathbf{x}_{t},\mathbf{g})\longrightarrow\{\mathbf{r}_{t,1},\mathbf{r}_{t,2},\dots\mathbf{r}_{t,N}\}$, that is used as context information to improve navigation.
-Depending on the use case, the use of observation $\mathbf{x}_{t}$ in the retrieval process may be optional — if not provided, retrieval can be done once per episode, as opposed to every time step.
+Depending on the use case, the use of observation $\mathbf{x}_{t}$ in the retrieval process may be optional (as in the static context in Section˜4) — if not provided, retrieval can be done once per episode, as opposed to every time step.
 […]
-$\begin{array}[]{lll}\tilde{\mathbf{x}}_{t}&=x(\mathbf{x}_{t})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// perceive}}\\ \tilde{\mathbf{g}}_{t}&=g(\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// compare obs+goal}}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{g}}_{t},l(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// update rec state}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// act}}\\ \end{array}$
-where $x$ and $g$ are trainable encoders, $h$ is the update function of a GRU [18] which maintains a hidden state $\mathbf{h}_{t}$ over time; $l$ is an embedding function, and $\pi$ is the policy.
+$\begin{array}[]{lll}\tilde{\mathbf{x}}_{t}&=x(\mathbf{x}_{t})&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// perceive}}\\ \tilde{\mathbf{g}}_{t}&=g(\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// compare obs+goal\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{g}}_{t},l(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// update recurrent state}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// act}}\\ \end{array}$
+where $x$ and $g$ are trainable encoders, $h$ is the update function of a GRU (Cho et al., 2014) which maintains a hidden state $\mathbf{h}_{t}$ over time; $l$ is an embedding function, and $\pi$ is the policy.
 For clarity we have omitted the equations of the gating functions.
 Then, the retrieval-augmented agent is defined as:
-$\begin{array}[]{lll}\tilde{\mathbf{x}}_{t}&=x(\mathbf{x}_{t})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// perceive}}\\ \tilde{\mathbf{g}}_{t}&=g(\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// compare obs+goal}}\\ \tilde{\mathbf{c}}_{t}&=c(\mathcal{R}(\mathbf{x}_{t},\mathbf{g}),\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// retrieve context}}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{g}}_{t},\tilde{\mathbf{c}}_{t},l(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// update rec state}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// act}}\\ \end{array}$
+$\begin{array}[]{lll}\tilde{\mathbf{x}}_{t}&=x(\mathbf{x}_{t})&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// perceive}}\\ \tilde{\mathbf{g}}_{t}&=g(\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// compare obs+goal}}\\ \tilde{\mathbf{c}}_{t}&=c(\mathcal{R}(\mathbf{x}_{t},\mathbf{g}),\mathbf{x}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// retrieve and encode context}}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{x}}_{t},\tilde{\mathbf{g}}_{t},\tilde{\mathbf{c}}_{t},l(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// update recurrent state}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.9609375,0.6953125,0.53125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9609375,0.6953125,0.53125}// act}}\\ \end{array}$
 where $c$ is an encoder which compresses the retrieved context into a compact form.
-It is trained end-to-end together with the agent, as described in Sec.
-3.1, while the retrieval process $\mathcal{R}$, described in Sec.
-4, leverages pre-trained models.
-In all tasks, the observation encoder $x$ is implemented as a half-width ResNet-18 (hwRN18) architecture [31].
-The goal $\mathbf{g}$, which is an image or a category, according to the navigation task, is compared to the observation $\mathbf{x}_{t}$ through a dedicated comparison function $g$ defined as
-$g(\mathbf{x}_{t},\mathbf{g})=\left\{\begin{array}[]{ll}d(\mathbf{x}_{t},\mathbf{g})&{\scriptsize\texttt{if ImageNav}}\\ [cl_{V}(\mathbf{x}_{t}),cl_{T}(\mathbf{g})]&{\scriptsize\texttt{if ObjectNav}}\\ \end{array}\right.$ (3)
-As in [6], for ImageNav we use a binocular visual encoder $d$ that compares the observation $\mathbf{x}_{t}$ with the goal image $\mathbf{g}$ and returns an embedding representing information about the goal’s direction, pose, and visibility.
-This encoder $d$ has been fine-tuned from the foundation model CroCo [74], solving as pre-text tasks relative-pose and visibility estimation [6].
-We use the weights available online.
-For ObjectNav where a goal category is specified, we use a CLIP encoder [59] providing a semantic comparison of the observed image, encoded as $cl_{V}(\mathbf{x}_{t})$, and the object category, encoded as $cl_{T}(\mathbf{g})$.
+It is trained end-to-end together with the agent, as described in Section˜3.1, while the retrieval process $\mathcal{R}$, described in Section˜4, leverages pre-trained models.
+The observation encoder $x$ is implemented as a half-width ResNet-18 (hwRN18) architecture (He et al., 2016).
+The image goal $\mathbf{g}$ is compared to the observation $\mathbf{x}_{t}$ through a function $g(\mathbf{x}_{t},\mathbf{g})$ implemented with a binocular visual encoder from Bono et al.
+(2024a) that returns an embedding representing information about the goal’s direction, pose, and visibility.
+This encoder was fine-tuned from the foundation model CroCo (Weinzaepfel et al., 2022), solving as pre-text tasks relative-pose and visibility estimation.
 Additionally, an image goal $\mathbf{g}^{\mathcal{R}}$ can be retrieved from $\mathcal{D}$ given the task goal $\mathbf{g}$, allowing the ImageNav model to address tasks in a zero-shot way.
 For example, an ImageNav agent can perform ObjectNav by retrieving a goal image $\mathbf{g}^{\mathcal{R}}$ from the database $\mathcal{D}$ given a goal category ([Uncaptioned image], [Uncaptioned image]).
-This can be integrated into the augmented agent above by adding a third case to the goal comparison function $g$,
-$g(\mathbf{x}_{t},\mathbf{g})=d(\mathbf{x}_{t},\mathbf{g}^{\mathcal{R}})\leavevmode\nobreak\ \leavevmode\nobreak\ {\scriptsize\texttt{if image goal isretrieved}}$ (4)
-We explore this use of retrieval for zero-shot task generalization in Sec. 5.2.
+In this case, the goal comparison function $g$ is computed between $\mathbf{x}_{t}$ and $\mathbf{g}^{\mathcal{R}}$ (instead of the original goal $\mathbf{g}$) as $g(\mathbf{x}_{t},\mathbf{g}^{\mathcal{R}})$.
+We explore this use of retrieval for zero-shot task generalization in Section˜5.2.
+Figure 2: We propose an architecture for Retrieval-Augmented Navigation architecture – RANa, that addresses all use cases of Fig.˜1b. Dashed arrows are optional depending on the navigation task.
 In summary, the agent receives at each timestep a new visual observation $\mathbf{x}_{t}$, which it encodes, compares with the goal $\mathbf{g}$, and enriches with additional context information retrieved from the database $\mathcal{D}$.
 This representation is then passed to a recurrent policy.
-See Fig.
-3 for a visual overview.
+See Fig.˜2 for a visual overview.
 3.1 Trainable context encoder $c$
 […]
-$\mathbf{e}_{t,n}=\leavevmode\nobreak\ \leavevmode\nobreak\ d(\mathbf{x}_{t},\mathbf{r}_{t,n})$ // ObjNav (5)
-$\mathbf{e}_{t,n}=[d(\mathbf{x}_{t},\mathbf{r}_{t,n})\ d(\mathbf{g},\mathbf{r}_{t,n})].$ // ObjNav or ImageNav (6)
+$\mathbf{e}_{t,n}=[g(\mathbf{x}_{t},\mathbf{r}_{t,n})\ g(\mathbf{g},\mathbf{r}_{t,n})].$ (3)
 At this point, we obtain a list of embeddings $\mathbf{e}_{t,n}$ that benefit from the rich inductive bias provided by the geometric FM, including information about pose and visibility of the context image $\mathbf{r}_{t,n}$ from the observed image $\mathbf{x}_{t}$ and goal image $\mathbf{g}$.
-We must now derive a single context embedding $\tilde{\mathbf{c}}_{t}=c(\mathcal{R}(\mathbf{x}_{t},\mathbf{g}),\mathbf{x}_{t},\mathbf{g})$ that provides useful information to the agent.
-We opt for two variants:
-Variant 1: Gumbel soft-max selector – The agent selects a single embedding from the list by predicting a distribution over items in the retrieved list followed by a Gumbel soft-max sampler [37].
-More precisely,
+The list is compressed into a single context embedding $\tilde{\mathbf{c}}_{t}=c(\mathcal{R}(\mathbf{x}_{t},\mathbf{g}),\mathbf{x}_{t},\mathbf{g})$ that provides useful information to the agent, and for this we explore two variants.
+Variant 1: Gumbel soft-max selector – The agent selects a single embedding from the list by predicting a distribution over items in the retrieved list followed by a Gumbel soft-max sampler (Jang et al., 2017):
 $\tilde{\mathbf{c}}_{t}\sim\sigma(\{\alpha_{n}\}_{n=1}^{N}),\quad\alpha_{n}=\texttt{Linear}(\mathbf{e}_{t,n})$ (4)
 where $\sigma$ is the soft-max distribution over items in the list.
-We perform the sampling and the computation of gradients using the Gumbel soft-max trick [37].
-We motivate this choice from the pre-training of the geometric FM $d$ in Eq.
-5 and Eq.
-6: the latent vector $z=d(a,b)$ has been pre-trained to provide overlap information between images $a$ and $b$, which we conjecture to be highly correlated with the relevance of the corresponding context item.
+We perform the sampling and the computation of gradients using the Gumbel soft-max trick (Jang et al., 2017).
+We motivate this choice from the pre-training of the geometric FM $g$ in Eq.˜3: the latent vector $z=g(a,b)$ has been pre-trained to provide overlap information between images $a$ and $b$, which we conjecture to be highly correlated with the relevance of the corresponding context item.
 Additionally, since the weights $\alpha_{i}$ are calculated independently, the model can be zero-shot applied to any context size at test time.
 Variant 2: Attention-based encoder – A transformer-based model is potentially capable of integrating multiple context tokens $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ into a single fused embedding, effectively changing and enriching the original embedding space.
-We use a transformer encoder network with 2 layers of 4-heads self-attention and predict the output by concatenating the context tokens and processing them with a 2-layer MLP.
+We use a transformer encoder layer with 2 layers of 4-heads self-attention, which receives as input the $N$ context tokens summed with positional embeddings, and predicts the context feature $\tilde{\mathbf{c}}_{t}$ by concatenating the output tokens and feeding them to a 2-layer MLP.
 4 Retrieval mechanisms for RANa
-In all tasks and use cases, the retrieval process consists of a nearest neighbor search on representations obtained with pre-trained models.
-Depending on the navigation task at hand, we either use DINOv2 [55] when we have a visual query or Open-CLIP [35] for textual queries.
-For zero-shot use-cases that require replacing the task goal $\mathbf{g}$ with a retrieved image goal $\mathbf{g}^{\mathcal{R}}$ ([Uncaptioned image], [Uncaptioned image]), we set $\mathbf{g}^{\mathcal{R}}$ as the nearest neighbor of $\mathbf{g}$ from the dataset $\mathcal{D}$.
-For use-cases that require building a context set to be passed to the encoder $c$ ([Uncaptioned image], [Uncaptioned image], [Uncaptioned image]), we investigate two strategies: (i) static context, where only the goal is used to build a fixed context kept throughout the episode, and (ii) dynamic context, which also depends on the current observation and is recomputed at each timestep.
+The retrieval process performs a nearest neighbor search on image representations obtained with DINOv2 (Oquab et al., 2024) resulting in a retrieved context.
+We investigate two context building strategies:
+  1. 1.
+    Static context, where only the goal is used to build a fixed context kept throughout the episode,
+  2. 2.
+    Dynamic context, which also depends on the current observation and is recomputed at each timestep.
 4.1 Static context
-We build the static context by ranking images from the dataset based on DINOv2 or CLIP feature similarity with the goal and selecting the top-$N$ nearest neighbors.
+[…]
 Diversity – providing multiple near-duplicates to the selector might not be very helpful.
-Moreover, in ObjectNav a scene may contain several instances of the target class and it is better to provide context images showing different instances.
-We therefore resort to Maximum Margin Relevance (MMR) [11] re-ranking to increase diversity.
-Assuming a relevance score vector $\omega$ (DINOv2 or CLIP similarities) providing a ranked shortlist (top-100 in our experiments), and a matrix $\Omega$ encoding the similarity for each pair of images in the ranked shortlist, MMR greedily selects at each step (rank) $i$ the element $r_{i}$ that maximizes the re-ranking criterion:
+We therefore resort to Maximum Margin Relevance (MMR) from Carbonell & Goldstein (1998) re-ranking to increase diversity.
+Assuming a relevance score vector $\omega$ (DINOv2 similarities) providing a ranked shortlist (top-100 in our experiments), and a matrix $\Omega$ encoding the similarity for each pair of images in the ranked shortlist, MMR greedily selects at each step (rank) $i$ the element $r_{i}$ that maximizes the re-ranking criterion:
 ${\text{MMR}(r_{i})=\beta\omega(r_{i})-\left(1-\beta\max_{r_{j}\in\mathcal{S}_{i-1}}\Omega(r_{i},r_{j})\right)},$
-where $\beta\in[0,1]$ is a mixture parameter (set to $0.5$ in all our experiments) and $\mathcal{S}_{i-1}$ is the set of images already selected. We then select the top-$N$ highest MMR.
-CLIP similarity normalization – While CLIP allows to easily compare any open-vocabulary object category with retrieval gallery images, it has been observed in the context of zero-shot image classification that directly ranking images by similarity to the category is suboptimal [58].
-To address this issue for ObjectNav, we leverage the fact that we often target a pre-defined set of classes $\mathcal{O}$ (e.g.
-6 in HM3D), and rescale the score of each image in $\mathcal{D}$ via softmax:
-${\omega_{\mathbf{o}}(\mathbf{x})=e^{cl_{S}(\mathbf{o},\mathbf{x})}\,/\,{\left(\sum_{\mathbf{o}^{\prime}\in\mathcal{O}}e^{cl_{S}({\mathbf{o}^{\prime},\mathbf{x}})}\right)}}$
-where $cl_{S}(\mathbf{o},\mathbf{x})$ denotes the CLIP similarity of image $\mathbf{x}$ and object category $\mathbf{o}$ and $\omega_{\mathbf{o}}(\mathbf{x})$ is its normalized similarity.
+[…]
 4.2 Dynamic context set
-It can be interesting to build the context using not only the goal but also the current observation, for example to find intermediate waypoints between the current observation and the goal.
-We propose to use an affinity graph based on similarities of images in the dataset $\mathcal{D}$ to find such waypoints.
-DINOv2-Graph – Given $\mathcal{D}$, we can build the affinity matrix $\Omega_{\mathcal{D}}$ containing the similarity between all pairs of images in $\mathcal{D}$ (we use DINOv2 [55] features).
-Using $\Omega_{\mathcal{D}}$, we can build a graph whose nodes are images $\mathbf{x}_{i}^{\mathcal{D}}\in\mathcal{D}$ and edges are weighted by similarity.
+Dynamic context exploits the current observation at each time step, and we leverage this to use the context set to provide intermediate waypoint images to the augmented agent.
+Similar to topological maps we do this by building a graph and calculating shortest paths.
+However, in stark contrast to classical topological maps, (i) the graph is built from the images of the retrieval dataset only and can be pre-computed, (ii) no pose information is required, we use visual similarities only, and (iii) waypoint images are recommendations only, exploited or not by the agent through its retrieval context.
+DINOv2-Graph – Given $\mathcal{D}$, we build the affinity matrix $\Omega_{\mathcal{D}}$ containing the similarity between all pairs of images in $\mathcal{D}$ using DINOv2 feature cosine similarity (Oquab et al., 2024).
+From $\Omega_{\mathcal{D}}$ we can derive a graph $\mathcal{G}=\{\mathcal{N},\mathcal{E}\}$, whose nodes $\mathcal{N}$ are images $\mathbf{x}_{i}^{\mathcal{D}}\in\mathcal{D}$ and edges $\mathcal{E}$ are weighted by similarity.
 Then, we find in $\mathcal{D}$ the most similar images $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ to the observation $\mathbf{x}_{t}$ and goal $\mathbf{g}$, respectively.
 Finally, we populate the context $\mathcal{R}(\mathbf{x}_{t},\mathbf{g})$ with $r_{1}^{\mathbf{x}_{t}}$ , $r_{1}^{\mathbf{g}}$ and images sampled on the shortest path in the graph between them.
-We provide more details about the graph and its relation to topological maps in the Supplementary.
+Fig.˜3 shows three examples of shortest paths on the DINOv2-Graph, in green.
+For comparison only, we also visualize the shortest paths obtained by the ground-truth (GT) “Pose Graph”, not available to the agent.
+The path on the DINOv2-Graph well approximates the GT path.
+We provide more details about the graph construction and its relation to topological maps in Appendix˜C.
+Figure 3: Shortest paths on DINOv2-Graph and Pose Graph - The most similar image $r_{1}^{\mathbf{x}_{t}}\in\mathcal{D}$ to the goal is highlighted in red and localized with a red star on the floorplan, while $r_{1}^{\mathbf{g}}\in\mathcal{D}$, in blue, is the most similar image to the current observation.
+Green crosses connects images along the shortest-path in the DINOv2-Graph (corresponding images in green) and purple crosses refer to poses on the Pose Graph, built using distances between GT camera poses, with corresponding images highlighted in purple at the bottom of the map.
+Retrieving the goal – independent of the creation of a retrieval context, static or dynamic, in zero-shot settings we also use retrieval to replace goal $\mathbf{g}$ with a retrieved image goal $\mathbf{g}^{\mathcal{R}}$ ([Uncaptioned image], [Uncaptioned image]), setting $\mathbf{g}^{\mathcal{R}}$ as the nearest neighbor of $\mathbf{g}$ from the dataset $\mathcal{D}$.
 5 Experimental results
-We train and evaluate our agents on the Habitat simulator and platform [64] according to the standard ImageNav, Instance-ImageNav and ObjectNav task definitions.
-Retrieval-augmented navigation benchmarks – There is no publicly available navigation benchmark that provides a database of observations that can be exploited during training and evaluation.
-We thus propose to derive ImageNav and ObjectNav benchmarks from standard Habitat tasks [64], modifying them as little as possible to enable training and evaluating retrieval-augmented agents.
-For ImageNav, we use the Gibson dataset [77], consisting of 72 train and 14 eval scenes, and for ObjectNav, we use 80 train and 20 eval scenes of the HM3DSem-v0.2 [60, 81].
-All agents use only one $112{\times}112$ RGB sensor.
-For each scene, unless otherwise stated, we generate a retrieval set $\mathcal{D}$ of $1,000$ FPV images by letting agents navigate the environment and store images in $\mathcal{D}$.
+We train and evaluate our agents on the Habitat simulator and platform (Savva et al., 2019) according to the standard ImageNav, Instance-ImageNav and ObjectNav task definitions.
+Retrieval-augmented navigation benchmarks – There is no publicly available benchmark that provides a database of observations that can be exploited during training and evaluation.
+We propose to derive our benchmarks from standard Habitat tasks (Savva et al., 2019), modifying them as little as possible to enable training and evaluating retrieval-augmented agents.
+We use the Gibson dataset (Xia et al., 2018), consisting of 72 train and 14 eval scenes, and for each scene, unless otherwise stated, we generate a retrieval set $\mathcal{D}$ of as little as 1,0001,0001 , 000 FPV images by letting agents navigate the environment and store images in $\mathcal{D}$.
 During training and evaluation, the retrieval-augmented agent can select dataset images from the scene it operates in.
-Further details on dataset creation are in Supplementary.
-Implementation details – For ImageNav ([Uncaptioned image], [Uncaptioned image]), we use the experimental setup for Gibson dataset [68] where the goal is a random view taken by the agent camera, as in, e.g.[48, 52, 6].
-We base our agents on the DEBiT architecture [6] by starting from the official codebase and weights provided by the authors [^2] https://github.com/naver/debit, and extend them as described in Sec.
-3.
-We consider two DEBiT backbones, base (DEBiT-b) with a binocular encoder with 55M parameters, and tiny (DEBiT-t), with a 17M-parameters encoder to allow faster execution for the ablation studies on Sec.
-5.4.
+Additional details on dataset creation are provided in Appendix˜A].
+Implementation details – We base our agents on the DEBiT architecture by starting from the official codebase and weights provided by the authors [^2] https://github.com/naver/debit, and extend them as described in Section˜3.
+We consider two DEBiT backbones, base (DEBiT-b) with a binocular encoder with 55M parameters, and tiny (DEBiT-t), with a 17M-parameters encoder, to allow faster execution of ablation studies.
 The corresponding retrieval-augmented models are RANa-b and RANa-t.
-In Sec.
-5.2 we use the same agent and weights in zero-shot fashion for Instance-ImageNav and ObjectNav when the task goal is replaced with a retrieved image goal ([Uncaptioned image], [Uncaptioned image]).
-For the ObjectNav base architecture with a textual goal ([Uncaptioned image], [Uncaptioned image]), we use the default experimental setup of HM3DSem-v0.2 [60, 10].
-We modify the RGB observations to be $112{\times}112$ instead of $480{\times}640$ as in the original ObjectNav definition, so we can load the pretrained weights of the hwRN18 visual encoder $x$ from DEBiT-t [6].
-We use the OpenCLIP [35] model trained on Laion-2B [65] [^3] https://github.com/mlfoundations/open_clip/.
-Model Retrieval Encoder SR SPL
-[Uncaptioned image] DEBiT-t [6] 79.9 51.4
-[Uncaptioned image] RANa-t DINOv2 Gumbel 83.0 52.3
+In Section˜5.2 we use the same agents and weights in zero-shot fashion for Instance-ImageNav and ObjectNav when the task goal is replaced with a retrieved image goal ([Uncaptioned image], [Uncaptioned image]).
+More details about the agent architecture are provided in Appendix˜B.
+Model training – All models, retrieval-augmented or not, are trained with PPO up to 200M steps.
+Unless specified otherwise, the geometric FM $g$ and encoders $x$ and $l$ are loaded from DEBiT and kept frozen, as is DINOv2.
+For retrieval-augmented agents, the context encoder $c$ and policy network $\pi$ are learned from scratch.
+Compared to DEBiT, the GRU memory of RANa receives the additional context input $\tilde{\mathbf{c}}_{t}$, therefore we expand the DEBiT GRU input matrix to reach the required dimension and finetune it.
+Further details and ablations of this choice are reported in Appendix˜D.
+The reward definition is inspired by PointNav (Chattopadhyay et al., 2021) and ImageNav (Bono et al., 2024a) rewards and given as:
+$r_{t}=\mathrm{K}\cdot\mathbf{1}_{\text{success}}-\Delta_{t}^{\mathrm{Geo}}-\lambda,$
+where $K{=}10$, $\Delta_{t}^{\mathrm{Geo}}$ is the increase in geodesic distance to the goal, and slack cost $\lambda{=}0.01$ encourages efficiency.
+Retrieval – We consider two context encoding approaches, Gumbel and Attention, and two retrieval mechanisms, static based on DINOv2, and dynamic based on DINOv2-Graph.
+The context size is $N{=}8$ in all experiments (see Appendix˜D for ablations).
+For dynamic variants requiring DINOv2 to be applied to each observation, we again leverage retrieval to speed up RL training significantly: during training, we replace observations $\mathbf{x}_{t}$ by their nearest dataset image, allowing to use pre-computed DINOv2 features.
+Metrics – Navigation performance is evaluated by success rate (SR), i.e., the percentage of episodes terminated within a distance of <1{<}1< 1m from the goal after the agent has called the STOP action, and SPL (Anderson et al., 2018), i.e., SR weighted by the optimality of the path:
+$\textit{SPL}=\frac{100}{N}\sum_{i=1}^{N}S_{i}\frac{\ell_{i}^{*}}{\max(\ell_{i},\ell_{i}^{*})},$
+where $S_{i}$ is a binary success indicator in episode $i$, $\ell_{i}$ is the agent path length and $\ell_{i}^{*}$ the shortest path length.
+5.1 Retrieval improves navigation
+Model Retrieval Context Enc. SR SPL
+[Uncaptioned image] DEBiT-t (Bono et al., 2024a) - - 79.9 51.4
+[Uncaptioned image] RANa-t static DINOv2 Gumbel 83.0 52.3
 [Uncaptioned image] RANa-t DINOv2 Attention 81.9 54.3
 [Uncaptioned image] RANa-t DINOv2-Graph Gumbel 82.0 51.5
-[Uncaptioned image] RANa-t DINOv2-Graph Attention 84.7 63.7
-[Uncaptioned image] DEBiT-b [6] 83.4 56.8
-[Uncaptioned image] RANa-b DINOv2 Gumbel 88.3 57.6
+[Uncaptioned image] RANa-t dynamic DINOv2-Graph Attention 84.7 63.7
+[Uncaptioned image] DEBiT-b (Bono et al., 2024a) - - 83.4 56.8
+[Uncaptioned image] RANa-b static DINOv2 Gumbel 88.3 57.6
 [Uncaptioned image] RANa-b DINOv2 Attention 84.3 58.3
 [Uncaptioned image] RANa-b DINOv2-Graph Gumbel 89.7 60.4
-[Uncaptioned image] RANa-b DINOv2-Graph Attention 90.7 71.8
-[Uncaptioned image] hwRN18+CLIP 42.9 23.0
-[Uncaptioned image] RANa-t OpenCLIP Gumbel 52.1 26.9
-[Uncaptioned image] RANa-t OpenCLIP Attention 49.7 26.0
-Table 2: Effect of retrieving context on ImageNav and ObjectNav. models finetuned for the same number of steps as corresponding RANa models for fairness of comparison.
-Model training – All models, retrieval-augmented or not, are trained with PPO up to 200M steps.
-Unless specified otherwise, the geometric FM $d$ and encoders $x$ and $l$ are loaded from DEBiT and kept frozen, as are OpenCLIP and DINOv2.
-For retrieval-augmented agents, the context encoder $c$ and policy network $\pi$ are learned from scratch.
-Compared to DEBiT and hwRN18+CLIP , the GRU memory of RANa receives the additional context input $\tilde{\mathbf{c}}_{t}$, therefore we expand the DEBiT/hwRN18+CLIP GRU input matrix to reach the required dimension and finetune it.
-Further details and ablations of this choice are reported in Supplementary.
-The reward definition is inspired by PointGoal [15] and ImageNav [6] rewards and given as $r_{t}=\mathrm{K}\cdot\mathbf{1}_{\text{success}}-\Delta_{t}^{\mathrm{Geo}}-\lambda,$ where $K{=}10$, $\Delta_{t}^{\mathrm{Geo}}$ is the increase in geodesic distance to the goal, and slack cost $\lambda{=}0.01$ encourages efficiency.
-Retrieval – We consider two context encoding approaches, Gumbel and Attention, and three retrieval mechanisms based on DINOv2 and DINOv2-graph for ImageNav, and OpenCLIP for ObjectNav tasks.
-Unless otherwise specified, we use a context size $N{=}8$.
-Results for ObjectNav use a retrieval dataset of size $10$k images (see Sec.
-5.4 and Supplementary for ablations).
-Metrics – Navigation performance is evaluated by the success rate (SR), i.e., the percentage of episodes terminated within a distance of ${<}1$m from the goal after the agent has called the STOP action, and SPL [2], i.e., SR weighted by the optimality of the path: $\textit{SPL}=\frac{100}{N}\sum_{i=1}^{N}S_{i}\frac{\ell_{i}^{*}}{\max(\ell_{i},\ell_{i}^{*})},$ where $S_{i}$ is a binary success indicator in episode $i$, $\ell_{i}$ is the agent path length and $\ell_{i}^{*}$ the shortest path length.
-5.1 Retrieval improves navigation
-Tab.
-2 shows results of the proposed RANa agents with retrieved context on ImageNav and ObjectNav tasks, which exhibit significant performance improvements across all settings compared to the DEBiT baseline and the ObjectNav baseline models [^4] To make baselines comparable, the loaded checkpoints continued training for the same additional 200M steps updating GRU and policy only..
-For static context retrieval (both using DINOv2 for ImageNav or OpenCLIP for ObjectNav) the Gumbel encoder shows better results, arguably because there is no inherent ordering in the context and the retrieved CroCo-based features contain strong information about the relative pose between context elements and observation.
+[Uncaptioned image] RANa-b dynamic DINOv2-Graph Attention 90.7 71.8
+Table 2: Effect of retrieving context, RANa-t and -b models.
+Table˜2 shows results of the proposed RANa agents with retrieved context on the ImageNav task, exhibiting significant performance improvements across all settings compared to the DEBiT baselines [^3] To make baselines comparable, loaded checkpoints continued training for the same additional 200M steps, updating GRU and policy only..
+For static context retrieval the Gumbel encoder shows better results, arguably because there is no inherent ordering in the context and the retrieved CroCo-based features contain strong information about the relative pose between context elements and observation.
+To simplify notation, from now on we label the combination of static context retrieval and Gumbel encoding static.
 The attention-based encoder works better for the dynamic graph-based context, where reasoning about the order of context elements becomes important.
-The graph-based context combined with attention obtains the best results, with a huge boost in SPL ($+14.2$ and $+12.3$ points) compared to the DEBiT-b and DEBiT-t baselines, respectively.
-For all ObjectNav runs, we found that finetuning $x$ while training on HM3D improved performance, as the visual encoder of DEBiT was only trained on Gibson observations.
-Moreover, since ImageNav and ObjectNav have different action spaces, we learn the action embedding $l$ from scratch.
-Model SR SPL Binocular encoder
-ZSON [47] 36.9 28.0 Observation finetuned, goal frozen
-OVRL [80] 54.2 27.0 Finetuned
-VC1-ViT-L [49] 81.6 - Finetuned
-OVRL-v2 [79] 82.0 58.7 Finetuned
-DEBiT-b no adapters [6] 83.0 55.6 Frozen
-DEBiT-L no adapters [6] 82.0 59.6 Frozen
-DEBiT-L+adapters[6] 94.0 71.7 Finetuned with adapters
-RANa-b Graph-Attn 90.7 71.8 Frozen (base DEBiT-b no adapters)
-Table 3: ImageNav state-of-the-art results. these models use additional GPS+Compass input providing exact localization.
-Comparison to ImageNav SotA – In Tab.
-3 we compare RANa ImageNav with state-of-the-art (SotA) models that typically finetune (with a navigation loss) visual encoders based on CLIP [47], ViT [80, 49, 79] and geometric FM [6] used for observation-goal comparison.
-RANa boosts the performance of DEBiT-b, which uses frozen binocular encoders, and achieves performance on par with SotA method DEBiT-Large [6] finetuned with adapters [17], which use an involving adaptation process to tune visual encoders to the task.
-In contrast, we achieve strong performance through (i) scalable context encoding and (ii) a generic frozen visual FM, enabling several zero-shot tasks (Sec.
-5.2).
-Comparison to ObjectNav SotATab.
-4 compares the results of RANa ObjectNav tiny model with competing approaches.
-Compared to methods of similar sensor settings (one FPV RGB camera), RANa-t outperforms previous SotA by $+9.7$ on SR and $+7.7$ on SPL.
-Neither ZSON [47] nor PSL [70] were trained on ObjectNav episodes, but on HM3D ImageNav episodes, which might give a small advantage to RANa.
-Hence we also compare RANa-t to methods that use extra sensors, such as depth camera, GPS+compass, or 6 camera sensors, as well as LLM-guided navigation.
-Here RANa-t shows competitive results, outperforming several strong methods such as ESC [86] and PixNav [10].
-Compared to the considerably more complex L3MVN [83], which leverages depth and pose to construct a semantic map and uses GPT-2 to guide exploration via object searching, RANa-t is competitive, being $-2.1$ below on SR but $+1.4$ above on SPL, meaning that the retrieved context guides our agent to more efficient routes to object goals.
-100 dataset images suffice to improve performanceTab.
-5 shows that the size of the retrieval database can be small, $100$ is enough for the considered Gibson scenes in ImageNav.
-Additional experiments for ObjectNav and zero-shot tasks are presented in Supplementary.
-Model Extra sensors w/ map w/ LLM SR SPL
-PixNav [10] Pano RGB 37.9 20.5
-ESC [86] Depth, odom 39.2 22.3
-L3MVN [83] Depth, odom 54.2 25.5
-CoW [27] Depth, odom 6.1 3.9
-GoW [27] Depth, odom 32.0 18.1
-ProcTHOR [19] Depth, odom 13.2 7.7
-ZSON [47] 25.5 12.6
-PSL [70] 42.4 19.2
-RANa-t (Ours) 52.1 26.9
-Table 4: Comparison to ObjectNav state-of-the-art methods.
-100 1k 10k 20k
+We label this graph-based context combined with attention encoding dynamic.
+It obtains the best results, with a huge boost in SPL (+15+15+ 15 and +12.3+12.3+ 12.3 points) compared to the DEBiT-b and DEBiT-t baselines, respectively.
+Model Odom SR SPL Binocular encoder
+OVRL (Yadav et al., 2023b) 54.2 27.0 Finetuned
+VC1-ViT-L (Majumdar et al., 2023) 81.6 - Finetuned
+OVRL-v2 (Yadav et al., 2023a) 82.0 58.7 Finetuned
+ZSEL (Al-Halah et al., 2022) 29.2 21.6 Obs. & policy frozen, goal from scratch
+ZSON (Majumdar et al., 2022) 36.9 28.0 Observation finetuned, goal frozen
+FGPrompt (Sun et al., 2024a) 92.3 68.5 Trained from scratch
+DEBiT-b no adapters (Bono et al., 2024a) 83.0 55.6 Frozen
+RANa-b dynamic 90.7 71.8 Frozen
+DEBiT-L w. adapters (Bono et al., 2024a) 94.0 71.7 Finetuned with adapters
+RANa-L w. adapters dynamic 94.5 76.2 Finetuned with adapters
+Table 3: ImageNav state-of-the-art results.
+this work sets the Habitat simulation parameter sliding=True, simplifying the task significantly and making it incomparable with other methods (cf.
+(Monaci et al., 2025)).
+Comparison to ImageNav SOTA – In Table˜3 we compare RANa with state-of-the-art (SOTA) ImageNav models.
+These typically finetune (with a navigation loss) visual encoders based on CLIP (Majumdar et al., 2022), ViT (Yadav et al., 2023b; Majumdar et al., 2023; Yadav et al., 2023a) and geometric FMs  (Bono et al., 2024a) used for observation-goal comparison.
+Not only RANa considerably boosts the performance of the strong DEBiT-b model, but it also enhance the SOTA performance of DEBiT-Large (Bono et al., 2024a) finetuned with adapters (Chen et al., 2023).
+While SR is only marginally improved, arguably because it is already very high (949494) and remaining failure cases are not solvable using context, path efficiency (expressed by SPL) increases significantly, from 71.771.771.7 to 76.276.276.2 points.
+[…]
+DEBiT-t 27M 79.3 50.0 13.2 - - - 1.7 1.1 10.4
+DEBiT-b 66M 83.0 55.6 25.2 - - - 1.7 1.1 22.4
+RANa-t static 32M 83.0 52.3 24.4 0.1 0.1 0.4 1.7 1.2 20.9
+RANa-t dynamic 34M 84.7 63.7 38.5 11.9 0.5 2.2 1.7 1.2 21.0
+Table 4: Inference runtime of different model components, in ms, timed on one Nvidia H100-80G GPU.
+Retrieval overhead is negligible, most runtime is spent on image feature computations (DINOv2, CroCo), which increase with retrieval because of additional operations needed to select and compare retrieved elements.
+Retrieval is efficientTable˜4 shows model size and runtime of different components of selected DEBiT and RANa models.
+Retrieval time is negligible and scales favorably with the dataset size thanks to efficient approximate nearest neighbor search libraries such as Faiss (Douze et al., 2024).
+Inference time is slightly higher for RANa because the encoder $c$ computes CroCo features between context images and observation and goal (while DEBiT only runs CroCo between observation and goal).
+Still, RANa-t is smaller and faster than DEBiT-b while achieving similar performance, suggesting that retrieval augmentation can be a frugal alternative to model scaling.
+The variant with dynamic context adds the compute time required to run DINOv2 on each observation, but its runtime remains in the same ballpark of other agents.
+100 dataset images suffice to improve performanceTable˜5 shows that the size of the retrieval database can be small, 100100100 images per scene suffice for the considered Gibson scenes, and performance plateaus at 111k images.
+Additional experiments for Instance-ImageNav and ObjectNav are presented in Appendix˜D.
+DB size 100 1k 10k 20k
 1k 81.7 (50.1) 82.7 (52.6) 82.1 (51.2) 82.8 (51.9)
 10k 80.6 (49.4) 83.0 (52.3) 79.8 (46.3) 82.2 (51.5)
-Table 5: [Uncaptioned image]  Impact of retrieval database size - SR (SPL) for ImageNav with retrieval using different database sizes, combinations of train and test conditions.
+Table 5: Impact of retrieval database size - SR (SPL) for RANa-t with static context using different database sizes, combinations of train and test conditions.
+[…]
+Model Retrieval Features SR SPL
+[Uncaptioned image] DEBiT-t - 79.9 51.4
+[Uncaptioned image] RANa-t static CroCo Geometric FM 83.0 52.3
+[Uncaptioned image] RANa-t static hwRN18 79.1 48.2
+Table 6: Impact of Geometric FM on context representation.
+Context is importantTable˜7 studies the behavior of RANa-t models when ablating the context content: (i) in domain: expected retrieved elements, and (ii) random: random images from the scene dataset.
+Providing random context images results in a performance drop of both RANa static and dynamic to baseline levels.
+This result supports two key properties of our approach: the context is useful and RANa learns to exploit it, and at same time RANa can filter out irrelevant context elements and is robust to misleading random retrieval.
+Model in domain random
+RANa-t static 83.0 (52.3) 77.0 (47.1)
+RANa-t dynamic 84.7 (63.7) 80.3 (50.2)
+Table 7: Impact of retrieved context - SR (SPL) for RANa-t models on different test conditions.
+Examples of retrieved contextFig.˜4 shows three types of context for $N=8$.
+Top 8 most similar dataset images to the goal in DINOv2 feature space (top) and top 8 MMR most similar images filtered by MMR (middle), are static context types, i.e.they are fixed along one episode and depend only on the goal image, shown on the right highlighted in red.
+DINOv2-Graph (bottom) is a dynamic context, recalculated at each step and depending on the observation (left, highlighted in blue).
+The top 8 most similar images to the goal (top) are very similar to each other as well, and might not be very useful.
+Filtering images using MMR increases diversity and performance (cf.
+Appendix˜D for a qualitative analysis).
+The dynamic context (bottom) is the most informative, as it provides some form of “waypoints” from the current location to the goal – indeed this approach achieves the best navigation performance, as shown in Tables˜2 and 3.
+Refer to captionStatic, top 8Static, top 8 MMRDynamic, DINOv2-Graph
+Figure 4: Context types - top 8 most similar dataset images to the goal (red, on the right), top 8 MMR, most similar images filtered by MMR, and dynamic, the only one depending on the observation (blue, left) as well.
 5.2 Retrieval allows zero-shot generalization
+ObjNav IIN
+Model Extra sensors SR SPL SR SPL
+ZSEL (Al-Halah et al., 2022) 11.3
+ZSON A (Majumdar et al., 2022) 31.3 12.0 16.1 8.4
+Mod-INN (Krantz et al., 2023) Depth, odom 56.1 23.3
+[Uncaptioned image] DEBiT-b 31.9 15.8 55.4 25.5
+[Uncaptioned image] RANa-b 36.6 17.9 58.7 26.8
+Table 8: Zero-shot navigation results using retrieved image goal $\mathbf{g}^{\mathcal{R}}$ for Gibson ObjectNav and HM3D Instance-ImageNav, and comparison to state-of-the-art.
+obtained with camera intrinsics and robot size of the original Instance-ImageNav task specification.
 By retrieving the image goal $\mathbf{g}^{\mathcal{R}}$ from the database, it is possible to apply any ImageNav model to the ObjectNav and Instance-ImageNav tasks.
-Tab.
-6 compares the performance of zero-shot ImageNav architectures in: (i) Gibson ObjectNav, a variant introduced in [1] for ImageNav agents, consisting of $1,000$ episodes over $5$ Gibson scenes and containing the 6 object categories of HM3D ObjectNav, and (ii) HM3D Instance-ImageNav with modified camera intrinsics and robot size to fit the ImageNav configuration.
-In both cases we use retrieval databases of $50,000$ images, since it is important to retrieve an image goal $\mathbf{g}^{\mathcal{R}}$ as close as possible to the goal, as shown in Sec.
-5.4.
+Table˜8 compares the performance of zero-shot architectures in: (i) Gibson ObjectNav, a variant introduced in Al-Halah et al.
+(2022) for ImageNav agents, consisting of 1,0001,0001 , 000 episodes over 555 Gibson scenes and containing the 6 object categories of HM3D ObjectNav (chair, bed, plant, sofa, tv, toilet), and (ii) HM3D Instance-ImageNav (Krantz et al., 2023) with modified camera intrinsics and robot size to fit the ImageNav configuration.
+In both cases we use retrieval databases of 50,00050,00050 , 000 images, as it is important to retrieve an image goal $\mathbf{g}^{\mathcal{R}}$ as close as possible to the real goal to avoid stopping too far from it.
+We ablate this parameter in Appendix˜D and show that RANa already achieves SOTA zero-shot Instance-ImageNav performance with as little as 5,0005,0005 , 000 retrieval images.
 In Instance-ImageNav we replace the image goal with the closest image in DINOv2 feature space from the retrieval set.
-In zero-shot ObjectNav we retrieve the top-9 closest images in CLIP space and re-rank them at each step to select as goal $\mathbf{g}^{\mathcal{R}}$ the element with the highest co-visibility with the current observation, and the remaining 8 images constitute the context.
+In zero-shot ObjectNav, where the goal is defined as an object category, we retrieve the top-9 closest dataset images in CLIP space (Ilharco et al., 2021) and re-rank them at each step to select as goal $\mathbf{g}^{\mathcal{R}}$ the element with the highest co-visibility with the current observation, and the remaining 8 images constitute the context.
 This is done using the frozen co-visibility auxiliary loss head used to train DEBiT models.
 This step is needed since in ObjectNav any instance of a target object is a valid goal.
-Retrieval enables the use of DEBiT-b for both tasks, achieving strong performance in line with best existing methods (for reference, DEBiT-b achieves $10$ SR and $3.1$ SPL on the original Instance-ImageNav task).
-RANa outperforms the feature-matching based method [42] which uses additional depth and odometry sensors and is the current SotA in zero-shot Instance-ImageNav, as well as ZSON (configuration A) [47] in zero-shot Gibson ObjectNav.
-ObjNav IIN
-Model Extra sensors SR SPL SR SPL
-ZER [1] 11.3
-ZSON A [47] 31.3 12.0 16.1 8.4
-Mod-INN [42] Depth, odom 56.1 23.3
-[Uncaptioned image] DEBiT-b 31.9 15.8 55.4 25.5
-[Uncaptioned image] RANa-b 36.6 17.9 58.7 26.8
-Table 6: Zero-shot navigation results using retrieved image goal $\mathbf{g}^{\mathcal{R}}$ for Gibson ObjectNav and HM3D Instance-ImageNav, and comparison to the state of the art.
-obtained with camera intrinsics and robot size of the original Instance-ImageNav task specification.
-Fig.
-4 shows an example of successful zero-shot Instance-ImageNavigation, where our RANa agent first correctly picks from the context the door near the goal ①, and then a view of the room where the target is ②, leading to success.
+Retrieval enables the use of DEBiT-b for both tasks, achieving strong performance in line with best existing methods (for reference, zero-shot DEBiT-b without retrieved goal image achieves 101010 SR and 3.13.13.1 SPL on Instance-ImageNav).
+RANa outperforms the feature-matching based method of Krantz et al.
+(2023) which uses additional depth and odometry sensors and is the current SOTA in zero-shot Instance-ImageNav, as well as ZSON (configuration A) of Majumdar et al.
+(2022) in zero-shot Gibson ObjectNav.
+Fig.˜5 shows an example of successful zero-shot Instance-ImageNavigation, where our RANa agent first correctly picks from the context the door near the goal ①, and then a view of the room where the target is ②, leading to success.
 In contrast, the DEBiT-b baseline without context does not see the target inside the room and passes by, exploring the wrong side of the house and failing to complete the episode.
-Figure 4: Zero-shot Instance-ImageNavigation episode by RANa and by the DEBiT-b baseline without context (top-right box).
-The green camera icon indicates the pose of the retrieved image goal $\mathbf{g}^{\mathcal{R}}$, the pink ones indicate the poses of selected context items.
+Figure 5: Zero-shot Instance-ImageNav episode by RANa-b and by DEBiT-b baseline (top-right box).
+The green camera indicates the pose of the retrieved image goal $\mathbf{g}^{\mathcal{R}}$, the pink ones indicate the poses of selected context items.
 5.3 Retrieval allows bridging the sim-to-real gap
-We evaluate our [Uncaptioned image] RANa-b agent on a small Instance-ImageNav dataset of $200$ episodes where goal images are pictures taken in a real office environment.
-We manually annotate the positions of $20$ real images in a simulated 3D reconstruction of the environment.
-We generate a retrieval set $\mathcal{D}$ of $10$k images from uniformly-sampled navigable poses, rendered in simulation using the agent configuration.
-Simulating navigation from $10$ random starting poses towards each of the $20$ real image goals, RANa reaches $46.5$ success, vs.
-$40.5$ of DEBiT-b (see inlet Table in Fig.
-5).
+We evaluate our [Uncaptioned image] RANa-b agent on a small Instance-ImageNav dataset of 200200200 episodes where goal images are taken in a real office environment.
+We manually annotate the positions of 202020 real images in a simulated 3D reconstruction of the environment.
+We generate a retrieval set $\mathcal{D}$ of 101010k images from uniformly-sampled navigable poses, rendered in simulation using the agent configuration.
+Simulating navigation from 101010 random starting poses towards each of the 202020 real image goals, RANa reaches 46.546.546.5 success, vs.
+40.540.540.5 of DEBiT-b (see inlet Table in Fig.˜6).
 The Figure illustrates a challenging success case (left), where the cactus viewing angle is far from any view the agent could capture, and (right) an interesting failure case, where the retrieval returned as goal $\mathbf{g}^{\mathcal{R}}$ an image of a different instance of the same microwave model at a different position.
-Interestingly, some retrieved context images depict the correct goal, and RANa appears to be able to exploit this information as it succeeds in $3/10$ episodes featuring this goal.
+Interestingly, some retrieved context images depict the correct goal, and RANa appears to be able to exploit this information as it succeeds in 3/103/103 / 10 episodes featuring this goal.
 Model SR SPL
@@
 Refer to caption RANa-b 46.5 17.0
-Figure 5: Navigation in simulation from real goal images, retrieving new goals: 2 episodes shown, 200 episodes evaluated.
-5.4 Further ablation studies
-All ablations were conducted on RANa-t models based on the DEBiT-t backbone to allow for faster training and evaluation. Additional ablations are reported in Supplementary.
-Model Retrieval Encoder Retr. Feature SR SPL
-[Uncaptioned image] DEBiT-t 79.9 51.4
-[Uncaptioned image] RANa-t DINOv2 Gumbel CroCo Geom. FM 83.0 52.3
-[Uncaptioned image] RANa-t DINOv2 Gumbel hwRN18 79.1 48.2
-Table 7: [Uncaptioned image]  Impact of Geometric FM on context representation.
-Geometric FM provides essential directional information – and plays a crucial role in the effectiveness of RANa.
-We validate this intuition by training with the same procedure as all other models a RANa-t ImageNav network where the encoders $d$ in Eq.
-6 are replaced by a hwRN18 (same as encoder $x$) that takes as $9{\times}112{\times}112$ input the channel-concatenated observation, goal and context, and outputs the context features $\mathbf{e}_{t,n}$, fed to the Gumbel selector.
-As shown in Tab.
-7, this model achieves $79.1$ SR, against $83.0$ SR of RANa-t: in this setting, the agent is not capable of exploiting context information and performance remain at DEBiT-t baseline level, validating the importance of the geometric pre-training done with the CroCo FM in [6].
-Retrieval is efficientTab.
-8 shows model size, retrieval ($t_{R}$) and inference ($t_{I}$) time for selected DEBiT and RANa models.
-Retrieval time is low (negligible in non-graph case), could even be further optimized by not retrieving at every step, and scales favorably with the dataset size thanks to efficient approximate nearest neighbor search libraries such as Faiss [22].
-Inference time is slightly higher for RANa compared to DEBiT because of encoder $c$ that computes CroCo features for context images, observation and goal.
-Still, RANa-t is smaller and faster than DEBiT-b while achieving similar performance, suggesting that retrieval augmentation can be a frugal alternative to model scaling.
-Context diversity is importantTab.
-9 evaluates the behavior of RANa-t ImageNav model when using different strategies to create the context, (i) rand: 8 random elements from the retrieval dataset, (ii) top 8: most similar database images to the goal in DINOv2 feature space, and (iii) top 8 MMR: most similar database images filtered by MMR as in Sec.
-4.1.
-Filtering the context with MMR is important, as the agent relies on diverse context items; removing MMR degrades performance.
-Other approaches have lower but more constant performance across settings.
-Similar experiments and behavior is presented for ObjectNav in Supplementary.
-Model Retrieval Nr Par $t_{R}(s)$ $t_{I}(s)$ SR SPL
-DEBiT-t 27M 0.016 79.3 50.0
-DEBiT-b 66M 0.025 83.0 55.6
-RANa-t DINO-Gumbel 32M 0.001 0.021 83.0 52.3
-RANa-t DINOGraph-Attn 34M 0.055 0.024 84.7 63.7
-Table 8: Retrieval and inference runtimes (Tesla-T4-16G GPU).
-rand top 8 top 8 MMR
-rand 81.9 (51.3) 80.3 (50.4) 81.0 (50.3)
-top 8 78.4 (49.0) 81.4 (48.1) 80.7 (51.1)
-top 8 MMR 77.0 (47.1) 73.7 (42.4) 83.0 (52.3)
-Table 9: [Uncaptioned image]  Impact of retrieval data selection - SR (SPL) for ImageNav with different context building strategies, combinations of train and test conditions.
+Figure 6: Navigation with real goal images, retrieving new goals: 2 episodes shown, 200 episodes evaluated.
+6 Extension to ObjectNav
+We extend the proposed approach to ObjectNav, where the goal is a textual label of an object category, by slightly modifying the RANa architecture.
+Since ObjectNav requires semantic understanding of the scene and common-sense reasoning, a large body of research builds semantic maps (Gadre et al., 2022; Chaplot et al., 2020a) and/or leverages an LLM to guide navigation based on a scene’s common sense semantics (Yu et al., 2023; Cai et al., 2024; Kuang et al., 2024; Yin et al., 2024).
+Few approaches address this task in a zero-shot fashion, training their agents without ObjectNav rewards and leveraging CLIP encoders for goal representation (Sun et al., 2024b; Majumdar et al., 2022; Gadre et al., 2023).
+Here we do not aim to achieve state-of-the-art results, as it is unlikely for our simple approach to compete with methods using LLMs and VLMs with billion of parameters.
+Instead we want to showcase the effectiveness and flexibility of our method.
+In order to do that, we simply modify few components of the RANa-t static architecture: to account for the textual nature of the goal we use the multi-modal FM OpenCLIP (Ilharco et al., 2021) to compute the goal representation $\tilde{\mathbf{g}}_{t}$ and the similarity measure used for retrieval.
+Besides, ObjectNav has two additional actions, LOOK UP and LOOK DOWN, that are added to the action space.
+We train and test this agent, as well as a baseline not augmented with retrieval (we refer to it as DEBiT-t ObjectNav), on the HM3DSem dataset (Cai et al., 2024; Ramakrishnan et al., 2021) with 6 object categories (chair, bed, plant, sofa, tv, toilet) using the standard Habitat ObjectNav task definition and following the same procedure described earlier.
+Table˜9 compares the results of RANa-t and DEBiT-t ObjectNav models with competing approaches.
+RANa can exploit retrieved context also in this setting and improves the DEBiT baseline SR by almost +10+10+ 10 points.
+Compared to methods with similar sensor settings (only one RGB camera), RANa-t outperforms previous SOTA by +9.7+9.7+ 9.7 on SR and +7.7+7.7+ 7.7 on SPL.
+We also compare RANa-t to methods that use extra sensors, such as depth camera, GPS+compass, or 6 camera sensors, as well as LLM-guided navigation.
+Here RANa-t shows competitive results, outperforming several strong baselines such as ESC (Zhou et al., 2023) and PixNav (Cai et al., 2024).
+Model Extra sensors w/ map w/ LLM SR SPL
+PixNav (Cai et al., 2024) Pano RGB 37.9 20.5
+ESC (Zhou et al., 2023) Depth, odom 39.2 22.3
+L3MVN (Yu et al., 2023) Depth, odom 54.2 25.5
+CoW (Gadre et al., 2022) Depth, odom 6.1 3.9
+GoW (Gadre et al., 2023) Depth, odom 32.0 18.1
+ProcTHOR (Deitke et al., 2022) Depth, odom 13.2 7.7
+OVRL-v2 (Yadav et al., 2023a) Odom 64.7 28.1
+ZSON (Majumdar et al., 2022) 25.5 12.6
+PSL (Sun et al., 2024b) 42.4 19.2
+DEBiT-t ObjectNav 42.9 23.0
+RANa-t ObjectNav 52.1 26.9
+Table 9: DEBiT-t and RANa-t ObjectNav compared to state-of-the-art methods.
 7 Conclusion
@@
 We leverage pre-trained foundation models to allow the agent to query and process contextual information that can be useful for solving its target task.
-Augmenting ImageNav and ObjectNav agents with a context retrieval module improves navigation performance in both tasks, despite the goal specifications by different modalities (an image in ImageNav , text in ObjectNav), which showcases the generality of the proposed method.
-Moreover, we show how to use retrieval across domains and modalities to apply existing agents to new tasks in a zero-shot fashion.
+Augmenting a SOTA ImageNav agent with a context retrieval module improves navigation performance while adding minimal overhead, which showcases the flexibility and effectiveness of the proposed method.
+Moreover, we demonstrate the approach for an ObjectNav agent, and we show how to use retrieval across domains and modalities to apply existing agents to new tasks in a zero-shot fashion.
 This work opens the possibility of letting navigation agents benefit from the large amount of unstructured data collected during the operation of robot fleets.
@@
 Appendix
-Appendix A Implementation details
-A.1 Retrieval dataset creation
-Both ImageNav and ObjectNav benchmarks use the Habitat simulator, platform and task definitions [64].
-For ImageNav, we use the Gibson dataset [77], consisting of 72 train and 14 eval scenes, and the ImageNav agent configuration: height 1.5m, radius 0.1m, $112{\times}112$ RGB camera with $90^{\circ}$ horizontal field-of-view (hfov) at 1.25m from the ground.
-For ObjectNav, we use 80 train and 20 eval scenes of the HM3DSem-v0.2 [60, 81] as in [10] and an agent of height 0.88m, radius 0.18m with $112{\times}112$ RGB sensor with $79^{\circ}$ hfov placed at 0.88m from the ground.
-For each scene we generate a retrieval dataset $\mathcal{D}$ usually containing $1,000$ FPV images as follows: (i) we randomly choose start and goal viewpoints such that the goal is navigable from the start, (ii) we make an agent follow the shortest path from start to goal, and (iii) after each action, we save the agent’s RGB observation as an image in $\mathcal{D}$.
+Appendix A Retrieval dataset creation
+We use the Habitat simulator, platform and task definitions from Savva et al.
+(2019).
+For ImageNav, we use the Gibson dataset (Xia et al., 2018), consisting of 72 train and 14 eval scenes, and the ImageNav agent configuration: height 1.5m, radius 0.1m, ${\times}112$ 112 RGB camera with $90^{\circ}$ horizontal field-of-view (hfov) at 1.25m from the ground.
+To train and test the ObjectNav variant in Section˜6, we use 80 train and 20 eval scenes of the HM3DSem-v0.2 (Ramakrishnan et al., 2021; Yadav et al., 2023c) as in Cai et al.
+(2024) and an agent of height 0.88m, radius 0.18m with ${\times}112$ 112 RGB sensor with $79^{\circ}$ hfov placed at 0.88m from the ground.
+For each scene, we generate a retrieval dataset $\mathcal{D}$ usually containing 1,0001,0001 , 000 FPV images as follows:
+  1. (i)
+    we randomly choose start and goal viewpoints such that the goal is navigable from the start,
+  2. (ii)
+    we make an agent follow the shortest path from start to goal, and
+  3. (iii)
+    after each action, we save the agent’s RGB observation as an image in $\mathcal{D}$.
 We repeat this process until the dataset contains the desired number of images.
 We randomly choose the distribution of starts and goals such that the agent roughly covers all navigable areas, and that the path between pairs of viewpoints are not too close to one another to simulate a prior offline exploration of the scene.
-A.2 RANa agent architecture
-In this section, we further explain some of the architecture choices made for our method.
+Appendix B RANa agent architecture
+In this section we detail some of the architecture choices made for our method.
 Observation encoderx processes the RGB input image $\mathbf{x}_{t}\in\mathbb{R}^{3\times 112\times 112}$.
-It is implemented with a half-width ResNet-18 [31] and generates an output feature $\tilde{\mathbf{x}}_{t}\in\mathbb{R}^{512}$.
+It is implemented with a half-width ResNet-18 (He et al., 2016) and generates an output feature $\tilde{\mathbf{x}}_{t}\in\mathbb{R}^{512}$.
 Goal comparison functiong compares the observation $\mathbf{x}_{t}$ to the goal $\mathbf{g}$.
-For ImageNav we use the frozen binocular encoder of DEBiT [6], based on the Geometric Foundation Model (FM) CroCo [74].
+For ImageNav we use the frozen binocular encoder of DEBiT (Bono et al., 2024a), based on the Geometric Foundation Model (FM) CroCo (Weinzaepfel et al., 2022).
 This encoder is a Siamese encoder $E$ applied to both $\mathbf{x}_{t}$ and $\mathbf{g}$, and a decoder $D$ which combines the output of the two encoders.
@@
 The output of the decoder $D$ is further compressed by a fully connected layer (frozen, from the DEBiT model) that projects the flattened output of $D$ into $\tilde{\mathbf{g}}_{t}\in\mathbb{R}^{3136}$.
-For details about the binocular encoder $g$ we refer to [6] and [74].
-In ObjectNav we use OpenCLIP [35] to encode the visual observation $\mathbf{x}_{t}$ and the textual goal $\mathbf{g}$.
+For details about the binocular encoder $g$ we refer to Bono et al.
+(2024a) and Weinzaepfel et al.
+(2022).
+In ObjectNav we use OpenCLIP (Ilharco et al., 2021) to encode the visual observation $\mathbf{x}_{t}$ and the textual goal $\mathbf{g}$.
 This process produces two features $cl_{V}(\mathbf{x}_{t})\in\mathbb{R}^{512}$ and $cl_{T}(\mathbf{g})\in\mathbb{R}^{512}$, which are concatenated to generate the output $\tilde{\mathbf{g}}_{t}\in\mathbb{R}^{1024}$.
-Action embeddingl encodes the previous action $\mathbf{a}_{t-1}$ into a $32$-dimensional feature vector $\tilde{\mathbf{a}}_{t}$ and it is taken from the DEBiT [6] model.
+Action embeddingl encodes the previous action $\mathbf{a}_{t-1}$ into a 323232-dimensional feature vector $\tilde{\mathbf{a}}_{t}$ and it is taken from the DEBiT (Bono et al., 2024a) model.
 For ImageNav, we load DEBiT action encoder and freeze it during training, while for ObjectNav, we trained it from scratch, since the two tasks do not share a common action space.
 Indeed, the task definition of ObjectNav requires two additional actions, LOOK UP and LOOK DOWN.
-Gumbel context encoderVariant 1 of the trainable context encoder c (Sec.
-3.1 ) selects one element $\tilde{\mathbf{c}}_{t}$ among the context features $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ based on a learned relevance estimate $\{\alpha_{n}\}_{n=1}^{N}$.
+Gumbel context encoderVariant 1 of the trainable context encoder c (Section˜3.1) selects one element $\tilde{\mathbf{c}}_{t}$ among the context features $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ based on a learned relevance estimate $\{\alpha_{n}\}_{n=1}^{N}$.
 The retrieved context images are compared to the goal image (if available) and to the observation by the same frozen binocular encoder of DEBiT described above, leading to the $N$ context features $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ each of dimension $\mathbb{R}^{D}$, where $D=6272$ for ImageNav or Instance-ImageNav (when goal image is available), and $D=3136$ for ObjectNav.
-A fully connected layer $\texttt{Linear}(D,1)$ (using a syntax inspired by PyTorch) estimates the relevance $\{\alpha_{n}\}_{n=1}^{N}$ for each context element and the item with largest relevance is selected using the Gumbel soft-max sampler [37].
-In the ImageNav or Instance-ImageNav cases, only the binocular feature representing the comparison between observation and context item is selected, so in all cases the context feature $\tilde{\mathbf{c}}_{t}\in\mathbb{R}^{3136}$.
-Attention-based context encoderVariant 2 of the trainable context encoder creates a context feature by computing the self-attention over the context elements $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ computed as described in the previous paragraph.
-In this case the context encoder c first reduces the dimensionality of each context feature to $\mathbb{R}^{128}$ through a fully connected layer $\texttt{Linear}(D,128)$ ($D=6272$ for ImageNav or Instance-ImageNav, and $D=3136$ for ObjectNav), then processes them with a standard two-layer transformer encoder network [72] with $4$ attention heads and feedforward dimension $256$.
-The $N$ context features are concatenated into a feature vector of size $\mathbb{R}^{N\times 128}$ and processed by a 2-layer $\texttt{MLP}(128,3136)$ with hidden dimension $128$ and output $3136$, such that the context encoder output $\tilde{\mathbf{c}}_{t}$ has the same dimension of that of the Gumbel context encoder.
-Recurrent memory – The flatten vectors $\tilde{\mathbf{x}}_{t}$, $\tilde{\mathbf{g}}_{t}$, $\tilde{\mathbf{a}}_{t}$ and $\tilde{\mathbf{c}}_{t}$ are concatenated and fed to a single-layer GRU [18] with hidden state $\mathbf{h}_{t}\in\mathbb{R}^{512}$.
+A fully connected layer $\texttt{Linear}(D,1)$ (using a syntax inspired by PyTorch) estimates the relevance $\{\alpha_{n}\}_{n=1}^{N}$ for each context element and the item with the largest relevance is selected using the Gumbel soft-max sampler (Jang et al., 2017).
+In the ImageNav or Instance-ImageNav cases, only the binocular feature representing the comparison between observation and context items is selected, so in all cases the context feature $\tilde{\mathbf{c}}_{t}\in\mathbb{R}^{3136}$.
+[…]
+Recurrent memory – The flatten vectors $\tilde{\mathbf{x}}_{t}$, $\tilde{\mathbf{g}}_{t}$, $\tilde{\mathbf{a}}_{t}$ and $\tilde{\mathbf{c}}_{t}$ are concatenated and fed to a single-layer GRU (Cho et al., 2014) with hidden state $\mathbf{h}_{t}\in\mathbb{R}^{512}$.
 This GRU is an “extended” version of the baseline agent’s GRU without context, which is finetuned.
@@
 Policy – The hidden state $\mathbf{h}_{t}$ is fed to a linear policy $\pi$ composed of a linear Actor head that generates a softmax distribution over actions and a linear Critic head that evaluates the current state.
-A.3 DINOv2-Graph
+Appendix C DINOv2-Graph
 In this section we provide details about the construction of the DINOv2-Graph context.
-The graph is built from the affinity matrix $\Omega_{\mathcal{D}}$ containing the similarity between DINOv2 [55] features of all pairs of images in $\mathcal{D}$, where the nodes are images $\mathbf{x}_{i}^{\mathcal{D}}\in\mathcal{D}$ and edges are weighted by similarity.
-We use Dijkstra’s algorithm [20] (from the SciPy python package) to select the shortest path between two nodes.
+The graph is built from the affinity matrix $\Omega_{\mathcal{D}}$ containing the similarity between DINOv2 (Oquab et al., 2024) features of all pairs of images in $\mathcal{D}$, where the nodes are images $\mathbf{x}_{i}^{\mathcal{D}}\in\mathcal{D}$ and edges are weighted by similarity.
+We use Dijkstra’s algorithm (Dijkstra, 1956) (from the SciPy python package) to select the shortest path between two nodes.
 Figure 7: Examples of shortest paths on DINOv2-Graph (SWG) and on Pose Graph (PS) - The most similar image $r_{1}^{\mathbf{x}_{t}}\in\mathcal{D}$ to the goal is highlighted in red and localized with a red star on the floorplan, while $r_{1}^{\mathbf{g}}\in\mathcal{D}$, in blue, is the most similar image to the current observation.
 Green crosses connects images along the shortest-path in the DINOv2-Graph (corresponding images are highlighted in green on top of the map) and purple crosses refer to poses on the Pose Graph, with corresponding images shown with purple contour at the bottom of the map.
-Selecting a path on this semantic graph (DINOv2-Graph) is not necessarily equivalent to a “physical path” (see examples in Fig.
-6).
-In order to enforce selection of edges between images that have high overlap probability (i.e.represent the same part of the scene) we made two further modifications to the above graph.
+Selecting a path on this semantic graph (DINOv2-Graph) is not necessarily equivalent to a “physical path” (see examples in Fig.˜7).
+In order to enforce selection of edges between images that have high overlap probability (i.e.
+represent the same part of the scene) we made two further modifications to the above graph.
 First, as we observed that a DINOv2 similarity score above $th=0.75$ between L2 normalized features reflects, in most cases, a good overlap between the two images, we only kept edges for which the similarity was above this threshold, removing all the others.
@@
 During both training and inference, at each step, the agent finds in $\mathcal{D}$ the most similar images $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ to the observation $\mathbf{x}_{t}$ and goal $\mathbf{g}$, respectively, and computes the shortest path between $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ on the DINOv2-Graph.
-Let $\mathcal{P}=\{r_{1}^{\mathbf{x}_{t}},r_{1},r_{2},...r_{n},r_{1}^{\mathbf{g}}\}$ be the set of nodes on the shortest path and $C$ the context size.
+Let $\mathcal{P}=\{r_{1}^{\mathbf{x}_{t}},r_{1},r_{2},...,r_{n},r_{1}^{\mathbf{g}}\}$ be the set of nodes on the shortest path and $C$ the context size.
 If $n+2>C$, we sample images randomly from $\mathcal{P}$ and update all images in the context.
@@
 If no path was found, $n=2$ hence we only update two elements of the context with $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$.
-SWG SBG DWG PG
-RANa-b (Gumbel) 89.7 (60.4) 89.6 (60.5) 88.7 (60.7) 89.7 (60.7)
-RANa-b (Attention) 89.8 (71.0) 89.6 (70.3) 90.7 (71.8) 91.9 (69.8)
-RANa-t (Attention) 84.7 (63.7) 83.5 (61.1) 86.8 (64.8) 87.4 (62.2)
-Table 10: [Uncaptioned image]  Impact of the graph structure on RANa with DINOv2-Graph - SR (SPL) for ImageNav with retrieval using different graph structures at test time.
-All RANa models were trained with Sparse Weighted Graph (SWG).
-In bolded best results the PG variant is not considered as it uses the camera poses of images in $\mathcal{D}$, we only add it for comparison.
-Analyzing the choice of the graph structure – We trained our models on the Gibson dataset where we had around $10$k images per scene in the retrieval dataset $\mathcal{D}$.
-We selected the checkpoint obtained at $130$M steps for all models.
+Analyzing the choice of the graph structure – We trained our models on the Gibson dataset where we had around 101010k images per scene in the retrieval dataset $\mathcal{D}$.
+We selected the checkpoint obtained at 130130130M steps for all models.
 During testing we used a set of $|\mathcal{D}|=1$k images per scene.
 In this section we study how the graph structure affects performance.
-The default model, used for the results reported in Tab.
-2 of the main paper, builds a weighted graph by removing edges with similarity below $th=0.75$ and uses $w_{ij}=\sqrt{(1-s_{ij})}$ as weights.
+The default model, used for the results reported in Table˜2 of the main paper, builds a weighted graph by removing edges with similarity below $th=0.75$ and uses $w_{ij}=\sqrt{(1-s_{ij})}$ as weights.
 We call this Sparse Weighted Graph (SWG).
@@
 (iii) Finally, we compare these results with a “pose graph”, where instead of using DINOv2 features, we use distances between camera poses to build the affinity map.
-In this case we consider a binary graph where we cut edges for which the distance between the camera poses is above $1$m.
+In this case we consider a binary graph where we cut edges for which the distance between the camera poses is above 111m.
 We denote the Pose Graph with PG.
-From the results shown in Tab.
-10 we can draw the following observations.
+SWG SBG DWG PG
+RANa-b (Gumbel) 89.7 (60.4) 89.6 (60.5) 88.7 (60.7) 89.7 (60.7)
+RANa-b (Attention) 89.8 (71.0) 89.6 (70.3) 90.7 (71.8) 91.9 (69.8)
+RANa-t (Attention) 84.7 (63.7) 83.5 (61.1) 86.8 (64.8) 87.4 (62.2)
+Table 10: [Uncaptioned image]  Impact of the graph structure on RANa with DINOv2-Graph - SR (SPL) for ImageNav with retrieval using different graph structures at test time: Sparse Weighted Graph (SWG), Sparse Binary Graph (SBG), Dense Weighted Graph (DWG) and Pose Graph (PG).
+All RANa models were trained with the same SWG graphs built on the training scenes.
+In bolded best results.
+Note that the PG variant is not considered as it uses the camera poses of images in $\mathcal{D}$; we only add it for comparison.
+From the results shown in Table˜10 we can draw the following observations.
 First, using a pose aware graph (PG) does not necessarily improve the accuracy over using DINOv2-Graph, and when it does, the improvement is rather small.
@@
 The second observation we can make is that the graph structure does not have a big impact on the model with a Gumbel selector, but a bigger one on the model with attention-based context encoder.
-This is not surprising as the Gumbel selects an image from the context and use it independently from other context elements, while Attention combines information from all images within the context, hence it is more important how we fill up the context.
+This is not surprising as the Gumbel selects an image from the context (path) and use it independently from other context elements, while Attention combines information from all images within the context (path), hence it is more important how we fill up the context.
 Finally, we observe that the Dense Weighted Graph (DWG) performs better than the Sparse Weighted Graph (SWG).
-We conjecture that the reason is that the model was trained with $\mathcal{D}$ of size $10$k images and we rarely had no connection between two requested nodes ($r_{1}^{\mathbf{x}_{t}}$ , $r_{1}^{\mathbf{g}}$), which occurs more often at test when only $1$k images are available.
-Using a denser graph (DWG) probably helps overcoming these cases.
-Similarities and differences with topological maps – The proposed graph representation is related to topological map structures [63, 4, 14, 76, 40, 69].
-However there are several critical differences between our model and traditional topological map based methods.
+We conjecture that the reason is that the model was trained with $\mathcal{D}$ of size 101010k images and we rarely had no connection between two requested nodes ($r_{1}^{\mathbf{x}_{t}}$ , $r_{1}^{\mathbf{g}}$), which occurs more often at test when only 111k images are available.
+Using a denser graph (DWG) appears to help overcoming these cases.
+Similarities and differences with topological maps – The proposed graph representation is related to topological map structures (Savinov et al., 2018; Beeching et al., 2020b; Chaplot et al., 2020c; Wiyatno et al., 2022; Kim et al., 2023; Sridhar et al., 2024).
+However, there are several critical differences between our model and traditional topological map based methods.
 First, we do not use the topological map to localize the agent, plan, and act to reach the “next waypoint”.
@@
 Additionally, our graph is connected based only on visual similarity rather than predictions from temporal distances.
-While it is debatable whether this approach is inherently superior, it offers a distinct advantage: we do not require trajectories to build the graph but can build it from a large number of isolated images.
-The main drawback is that we rely on the assumption that visual proximity is a proxy for spatial proximity, which is not always true (see Fig.
-6).
-Future work could attempt to learn to take into account the uncertainty in this assumption through data-driven planning, as in [4].
-Appendix B Additional ablation experiments
-In this section, we provide additional ablation studies that could not fit into the main document.
-As for the experiments in Sec.
-5.4 of the main paper, all experiments are conducted on RANa-t models with Gumbel selector as context encoder $c$.
-4 8 12 16 20 24
+While it is debatable whether this approach is inherently superior, it offers a distinct advantage: we do not require trajectories to build the graph but we can also build it from a large number of isolated images.
+The main drawback of our method is that we rely on the assumption that visual proximity is a proxy for spatial proximity, which is not always true (see Fig.˜7).
+Future work could attempt to learn to take into account the uncertainty in this assumption through data-driven planning, as in Beeching et al.
+(2020b).
+Appendix D Ablation experiments
+In this section, we provide additional ablation studies that could not fit into the main document. As for the ablation experiments in the main paper, all experiments are conducted on RANa-t models.
+Impact of context size – is studied in Table˜11 for ImageNav ([Uncaptioned image]) and in Table˜12 for ObjectNav ([Uncaptioned image]) with static context.
+One interesting advantage of the Gumbel selector is that it allows for context of different sizes during evaluation, independent of the context size used during training.
+With that said, this parameter has a relatively minor impact on performance, with larger training context sizes showing positive performance trends.
+The size $N=8$ is a choice that works well both for ImageNav and ObjectNav and balances navigation performance with execution time.
+For ObjectNav, we observe a more pronounced performance increase for larger context sizes during evaluation, so we use $N=8$ during training and $N=12$ during evaluation.
+context size 4 8 12 16 20 24
 4 80.8 (52.3) 80.5 (51.6) 80.4 (52.7) 81.0 (52.3) 80.9 (52.0) 81.7 (54.0)
@@
 Table 11: [Uncaptioned image]  Impact of context size for RANa-t - SR (SPL) for ImageNav with retrieval over different context sizes, combinations of train and test conditions.
-4 8 12 16 20 24
+context size 4 8 12 16 20 24
 4 51.6 (26.1) 50.8 (25.7) 49.8 (25.3) 49.8 (25.3) 49.0 (24.7) 50.6 (25.8)
@@
 Table 12: [Uncaptioned image]  Impact of context size for RANa-t - SR (SPL) for ObjectNav with retrieval over different context sizes, combinations of train and test conditions.
-Impact of context size – is studied in Tab.
-11 for ImageNav ([Uncaptioned image]) and in Tab.
-12 for ObjectNav ([Uncaptioned image]).
-One interesting advantage of the Gumbel selector is that it allows for context of different sizes during evaluation, independent of the context size used during training.
-With that said, this parameter has a relatively minor impact on performance, with larger training context sizes showing positive performance trends.
-The size $N=8$ is a choice that works well both for ImageNav and ObjectNav and balances navigation performance with execution time.
-For ObjectNav, we observe a more pronounced performance increase for larger context sizes during evaluation, so we use $N=8$ during training and $N=12$ during evaluation.
-[…]
-Model Context SR SPL
-RANa-t top 8 83.0 52.3
-RANa-t top 1 79.6 48.7
-Figure 7: Refer to caption  Histogram of the rank of the context element selected by RANa-t with Gumbel context encoder on the ImageNav Gibson eval set.
-Impact of retrieval database size – on zero-shot Instance-ImageNav ([Uncaptioned image]) is shown on Tab.
-13.
+Impact of context diversity for Gumbel context encoderTable˜13 evaluates the behavior of RANa-t ImageNav with static context, when using different strategies to create the context, (i) top 8 MMR: most similar database images filtered by MMR as in Section˜4.1, (ii) top 8: most similar database images to the goal in DINOv2 feature space, and (iii) top 1: the closest element from the retrieval dataset.
+Filtering the context with MMR is important, as the agent appears to rely on diverse context items; removing MMR degrades performance.
+top 1 always uses the best element from the context, which is very similar to the goal image.
+In this case performance go back to baseline levels, confirming the importance of diversity in the retrieved elements.
+top 8 MMR top 8 top 1
+top 8 MMR 83.0 (52.3) 81.5 (51.0) 79.6 (48.7)
+top 8 80.7 (51.1) 81.4 (48.1) 80.2 (51.2)
+Table 13: Impact of MMR filtering on retrieved context - SR (SPL) for ImageNav with and without MMR, combinations of train and test conditions ( in-domain performance shaded in pink).
+Context diversity from MMR is important and delivers the best performance (SR=83.0).
+When the model trained with MMR is tested without it, performance drops.
+Impact of retrieval database size – on zero-shot Instance-ImageNav ([Uncaptioned image]) is shown on Table˜14.
 Here the goal image $\mathbf{g}^{\mathcal{R}}$ is retrieved from the database, and thus retrieval accuracy is crucial for success and avoid terminating an episode too far from the actual goal.
-Tab.
-13 shows this phenomenon, with zero-shot Instance-ImageNav performance clearly increasing with retrieval dataset size.
-1k 5k 10k 20k 50k 100k
-47.6 (20.2) 56.2 (24.4) 56.4 (23.8) 57.4 (25.0) 58.7 (26.8) 59.1 (25.3)
+Table˜14 shows this phenomenon, with zero-shot Instance-ImageNav performance increasing with retrieval dataset size.
+However, please note that RANa achieves navigation performance in line with SOTA using a retrieval dataset containing only 5,0005,0005 , 000 images.
+DB size 1k 5k 10k 20k 50k 100k
+SR (SPL) 47.6 (20.2) 56.2 (24.4) 56.4 (23.8) 57.4 (25.0) 58.7 (26.8) 59.1 (25.3)
 Table 14: [Uncaptioned image] Impact of retrieval database size in zero-shot - SR (SPL) in Instance-ImageNavigation increases with dataset size.
-Analysis of “oracle” context – for ImageNav ([Uncaptioned image]) is carried out in Tab.
-14.
+Analysis of “oracle” context – for ImageNav ([Uncaptioned image]) is carried out in Table˜15.
 In this experiment we study which information is most useful and can be exploited by RANa, by building contexts using privileged information in simulation – therefore in these cases the retrieval dataset $\mathcal{D}$ is not used.
 Oracle 1 captures a panoramic view of the goal location and is created by collecting $N=8$ context images at the goal position while rotating the camera view by $45^{\circ}$ clock-wise for each element.
-This context represents a soft upper bound for the static context in Sec.
-4.1 , where a fixed set of items is chosen per episode based on their similarity with the goal image.
+This context represents a soft upper bound for the static context in Section˜4.1, where a fixed set of items is chosen per episode based on their similarity with the goal image.
 Oracle 2 context is richer and it is generated by computing the shortest path from the agent position to the goal at each step and collecting views along this path.
-This is a soft upper bound for the DINOv2-Graph context (Sec.
-4.2 ), which is dynamically generated at each step on the context graph.
-Tab.
-14 shows that, as expected, both types of oracle context can provide large gains, especially in SPL, i.e.navigation efficiency.
-The Table suggests that there is little room for improvement in success (SR) using static goal context, while more structured information, e.g.paths, derived from data in $\mathcal{D}$ can improve navigation considerably.
+This is a soft upper bound for the DINOv2-Graph dynamic context (Section˜4.2), which is dynamically generated at each step on the context graph.
+Table˜15 shows that, as expected, both types of oracle context can provide large gains, especially in SPL, i.e.
+navigation efficiency.
+It suggests that there is little room for improvement in success (SR) using static goal context, while more structured information, e.g.
+paths, derived from data in $\mathcal{D}$ can improve navigation considerably.
 Context Type SR SPL
-RANa-t goal static 83.0 52.3
+DINOv2 goal static 83.0 52.3
 Panorama (oracle 1) static 83.4 71.6
-RANa-t DINOv2-graph dynamic 84.7 63.7
+DINOv2-graph dynamic 84.7 63.7
 Shortest path (oracle 2) dynamic 98.3 87.9
-Table 14: [Uncaptioned image]  Navigation with “oracle” context - SR (SPL) for ImageNav using privileged information to build oracle context.
+Table 15: [Uncaptioned image]  Navigation with “oracle” context - SR (SPL) for RANa-t using privileged information to build oracle context.
 Impact of using pre-trained GRU weights – As mentioned earlier all RANa agents use a GRU finetuned from the baseline agent’s GRU.
-We evaluate the impact of this choice for the ImageNav ([Uncaptioned image]) model by training the RANa-t GRU from scratch while keeping all other settings fixed.
-Tab.
-15 shows that finetuning the GRU form pretrained weights – those of DEBiT [6] in this case – improves navigation performance.
+We evaluate the impact of this choice by training the RANa-t static GRU from scratch while keeping all other settings fixed.
+Table˜16 shows that finetuning the GRU form pretrained weights – those of DEBiT (Bono et al., 2024a) in this case – improves navigation performance.
 We also observe faster convergence during training, which might in part explain the performance increase.
@@
 GRU from scratch 81.4 51.0
-Table 15: [Uncaptioned image]  Impact of GRU finetuning - SR (SPL) for ImageNav with GRU finetuned and trained from scratch.
-Context selection criteria – While in RANa ObjectNav the context selection is always based on the observation, as the goal image is not available, in RANa ImageNav model we have both observation and goal images.
-All ImageNav experiments reported in the paper use a feature vector for each context element which is obtained by concatenating two features: (i) the output of the Geometric FM $d$ comparing context element and observation, and (ii) the output of $d$ comparing context element and goal.
-We evaluate the impact of this choice by training a RANa-t model which selects context elements based only on feature (i) which compares context elements to the observation.
-The performance of this variant, compared to the baseline which uses both observation and goal, is displayed in Tab.
-16, which shows that information provided by comparing the context to the goal is useful and can be exploited by RANa to navigate more successfully and efficiently to the goal.
+Table 16: [Uncaptioned image]  Impact of GRU finetuning - SR (SPL) for RANa-t with static context and GRU finetuned or trained from scratch.
+Context selection criteria – All ImageNav experiments reported in the paper use a feature vector for each context element which is obtained by concatenating two features: (i) the output of the Geometric FM $g$ comparing context element and observation, and (ii) the output of $g$ comparing context element and goal.
+We evaluate the impact of this choice by training a RANa-t static model which selects context elements based only on feature (i), which compares context elements to the observation.
+The performance of this variant and that of the baseline which uses both observation and goal, is displayed in Table˜17.
+It shows that information provided by comparing the context to the goal is useful and can be exploited by RANa to navigate more successfully and efficiently to the goal.
 Model SR SPL
@@
 context selection based on obs only 81.0 50.3
-Table 16: [Uncaptioned image]  Impact of context selection - SR (SPL) for ImageNav with context elements selected based on comparisons with only the observation, or both goal and observation.
-Context building for ObjectNavTab.
-17 shows the behavior of RANa-t ObjectNav ([Uncaptioned image]) when adopting different context building approaches, (i) rand: 8 random elements from the retrieval dataset, (ii) top 8 $cl_{S}$: most similar database images to the object goal encoder in CLIP feature space using CLIP similarity as metric, and (iii) top 8 $\omega_{\mathbf{o}}$: most similar database images using the score softmax normalization described in Sec.
-(4.1 ) of the main paper.
-The observed behavior is similar to ImageNav (Sec.
-5.4 ).
+Table 17: [Uncaptioned image]  Impact of context selection - SR (SPL) for RANa-t with static context and elements selected based on comparisons with only the observation, or both goal and observation.
+Context building for ObjectNavTable˜18 shows the behavior of RANa-t ObjectNav ([Uncaptioned image]) when adopting different context building approaches, (i) rand: 8 random elements from the retrieval dataset, (ii) top 8 $cl_{S}$: most similar database images to the object goal encoder in CLIP feature space using CLIP similarity as metric, and (iii) top 8 $\omega_{\mathbf{o}}$: most similar database images using CLIP score softmax normalization.
+While CLIP allows to easily compare any open-vocabulary object category with retrieval gallery images, it has been observed in the context of zero-shot image classification that directly ranking images by similarity to the category is suboptimal (Qian et al., 2023).
+To address this issue for ObjectNav, we leverage the fact that we often target a pre-defined set of classes $\mathcal{O}$ (e.g.
+6 in HM3D), and rescale the score of each image in $\mathcal{D}$ via softmax:
+${\omega_{\mathbf{o}}(\mathbf{x})=e^{cl_{S}(\mathbf{o},\mathbf{x})}\,/\,{\left(\sum_{\mathbf{o}^{\prime}\in\mathcal{O}}e^{cl_{S}({\mathbf{o}^{\prime},\mathbf{x}})}\right)}}$
+where $cl_{S}(\mathbf{o},\mathbf{x})$ denotes the CLIP similarity of image $\mathbf{x}$ and object category $\mathbf{o}$ and $\omega_{\mathbf{o}}(\mathbf{x})$ is its normalized similarity.
+The observed behavior is similar to ImageNav.
 First, a RANa-t model trained with a context with randomly selected images from the scene gallery performs worse than a model trained with either top 8 images.
-Nevertheless, this model still outperforms our baseline trained without retrieval (SR=$42.9$ and SPL=$23.0$ for the non-retrieval baseline) and is more robust to random context images during evaluation.
-For RANa-t models trained with retrieval top 8 context, we observe the improved performance of using softmax normalized scores.
+Nevertheless, this model still outperforms our baseline trained without retrieval (SR=42.942.942.9 and SPL=23.023.023.0 for the non-retrieval baseline) and is more robust to random context images during evaluation.
+For RANa-t models trained with retrieval top 8 context, we observe improved performance when using softmax normalized scores.
 rand top 8 $cl_{S}$ top 8 $\omega_{\mathbf{o}}$
@@
 top 8 $\omega_{\mathbf{o}}$ 44.4 (21.4) 51.3 (25.9) 51.9 (26.4)
-Table 17: [Uncaptioned image]  Impact of retrieval data selection - SR (SPL) for ObjectNav with different context building strategies, combinations of train and test conditions.
-Impact of retrieval dataset size for ObjectNav – is shown in Tab.
-18.
-Results demonstrate that unlike ImageNav (Tab.
-5 ), ObjectNav seems to need larger retrieval datasets both at training and at testing.
-Performance degrades significantly when the dataset $\mathcal{D}$ at training only has $1,000$ images.
-At test, using larger retrieval datasets leads to better performance, with metrics degrading significantly with $\mathcal{D}$ containing $100$ items.
-100 1k 10k 50k
+Table 18: [Uncaptioned image]  Impact of retrieval data selection - SR (SPL) for RANa-t ObjectNav with different context building strategies, combinations of train and test conditions.
+Impact of retrieval dataset size for ObjectNav – is shown in Table˜19.
+Results demonstrate that unlike ImageNav (Table˜5), ObjectNav seems to need larger retrieval datasets both at training and at testing.
+Performance degrades significantly when the dataset $\mathcal{D}$ at training only has 1,0001,0001 , 000 images.
+At test, using larger retrieval datasets leads to better performance, with metrics degrading significantly with $\mathcal{D}$ containing 100100100 items.
+DB size 100 1k 10k 50k
 1k 43.8 (21.7) 45.3 (22.4) 46.8 (23.2) 47.1 (23.2)
 10k 46.6 (24.9) 50.3 (26.4) 52.1 (26.9) 51.8 (27.2)
-Table 18: [Uncaptioned image]  Impact of retrieval database size - SR (SPL) for ObjectNav with retrieval using different database sizes, combinations of train and test conditions.
-Impact of CLIP variants for ObjectNav – is evaluated in Tab.
-19.
-In this experiment, we build retrieval datasets $\mathcal{D}$ of size $10$k images for each of the 100 HM3DSem-v0.2 scenes used for ObjectNav (80 train, 20 test).
-For each of the 6 target object categories (chair, bed, plant, toilet, tv monitor, sofa) and for each scene, we do a search in the retrieval dataset $\mathcal{D}$ by querying the CLIP-encoded goal category (i.e.the CLIP text encoder for the sentence “a picture of a category inside a house”) and comparing it with CLIP encoded dataset images of the scene, providing top 1, top 8 and top 20 success percentages.
-A retrieval is considered successful if the retrieved image (whose ground-truth pose is known), is closer to one of the viewpoints of a object instance of the correct category than the success threshold of the ObjectNav task, i.e.1111m.
-The Table shows that OpenCLIP performs better than alternative CLIP models, namely OpenAI’s original CLIP weights [59] and SigLIP [85], all with the softmax normalization proposed in Sec.
-4.1 being used.
-As observed earlier, ranking images by CLIP similarity to the category does not perform well [58]; since the number of target classes is known (6 in all our experiments), rescaling the score of each image in $\mathcal{D}$ via softmax significantly boosts performance.
-Vision-language model softmax (Sec. 4.1 ) top 1 top 8 top 20
-OpenCLIP [35] 40.6 76.0 88.5
-OpenCLIP [35] 78.1 95.8 97.9
-CLIP [59] 67.7 88.5 90.6
-SigLIP [85] 70.8 91.7 94.8
-Table 19: Retrieval performance of CLIP variants in the context of ObjectNav ([Uncaptioned image]).
-Retrieval is considered success if the retrieved image is of the correct category and if it is closer than $1$m to the target object.
-Appendix C Qualitative analysis
-C.1 Examples of retrieved context
-In this section we provide visual examples of the different context types used by our RANa model.
-Fig.
-8 shows three types of context used in this work for RANa ImageNav ([Uncaptioned image]), for $N=8$: top 8 most similar dataset images to the image goal in DINOv2 feature space (top) and top 8 MMR most similar images filtered by Maximum Margin Relevance (MMR) (middle), are static context types, i.e.they are fixed along one episode and depend only on the goal image, shown on the right.
-DINOv2-Graph (bottom) is a dynamic context that is recalculated at each step and depends on the observation (left) as well.
-This example clearly shows that the most similar images to the goal are very similar to each other as well, and might not be very useful to the RANa agent.
-Filtering images using MMR increases the diversity of the context items, and this, as shown in Tab.
-9 in the main text, increases RANa performance.
-For navigation purposes, the DINOv2-Graph context seems the most informative, as it provides waypoints from the current location to the goal – indeed this context provides the best navigation performance for RANa, see Tab.
-2 .
-Figure 8: Refer to caption  Context types for ImageNav - From top to bottom: top 8 most similar dataset images to the image goal (right), top 8 MMR most similar images filtered by MMR, and DINOv2-Graph, the only one depending on the observation (left) as well.
-Fig.
-9 displays examples of context for RANa ObjectNav ([Uncaptioned image]) using similarity in CLIP feature space for the six different goal categories, chair, bed, plant,sofa, tv monitor and toilet.
-As discussed in the paper, results accurately represent instances of the target object present in the scene, with the exception of the tv monitor class where we observe some errors, especially in scenes where monitors are absent.
+Table 19: [Uncaptioned image]  Impact of retrieval database size - SR (SPL) for RANa-t ObjectNav with retrieval using different database sizes, combinations of train and test conditions.
+Impact of CLIP variants for ObjectNav – is evaluated in Table˜20.
+In this experiment, we build retrieval datasets $\mathcal{D}$ of size 101010k images for each of the 100 HM3DSem-v0.2 scenes used for ObjectNav (80 train, 20 test).
+For each of the 6 target object categories (chair, bed, plant, toilet, tv monitor, sofa) and for each scene, we do a search in the retrieval dataset $\mathcal{D}$ by querying the CLIP-encoded goal category (i.e.
+the CLIP text encoder for the sentence “a picture of a category inside a house”) and comparing it with CLIP encoded dataset images of the scene, providing top 1, top 8 and top 20 success percentages.
+A retrieval is considered successful if the retrieved image (whose ground-truth pose is known), is closer to one of the viewpoints of a object instance of the correct category than the success threshold of the ObjectNav task, i.e.
+111m.
+The Table shows that OpenCLIP performs better than alternative CLIP models, namely OpenAI’s original CLIP weights (Radford et al., 2021) and SigLIP (Zhai et al., 2023), all with the softmax normalization being used.
+As observed earlier, ranking images by CLIP similarity to the category does not perform well (Qian et al., 2023); since the number of target classes is known (6 in all our experiments), rescaling the score of each image in $\mathcal{D}$ via softmax significantly boosts performance.
+Vision-language model softmax top 1 top 8 top 20
+OpenCLIP (Ilharco et al., 2021) 40.6 76.0 88.5
+OpenCLIP (Ilharco et al., 2021) 78.1 95.8 97.9
+CLIP (Radford et al., 2021) 67.7 88.5 90.6
+SigLIP (Zhai et al., 2023) 70.8 91.7 94.8
+Table 20: Retrieval performance of CLIP variants. Retrieval is considered success if the retrieved image is of the correct category and if it is closer than 111m to the target object.
+Appendix E Qualitative analysis
+E.1 Examples of retrieved context for ObjectNav
+Fig.˜8 displays examples of context for RANa ObjectNav ([Uncaptioned image]) using similarity in CLIP feature space for the six different goal categories, chair, bed, plant, sofa, tv monitor and toilet.
+Results accurately represent instances of the target object present in the scene, with the exception of the tv monitor class where we observe some errors, especially in scenes where monitors are absent.
 Figure 8: Refer to caption  Examples of context for different ObjectNav categories.
-C.2 Navigation examples
+E.2 Navigation examples
 Figure 9: Refer to caption  RANa-b successfully navigates to challenging goals in two ImageNav episodes - The agent observation is on the left, next to it is the goal image highlighted in green, then there are $N=8$ context elements, with the one selected by the RANa agent at the current time step highlighted in red, and on the right the agent path drawn on the floorplan of the scene, with the goal position indicated by a red square.
 At the bottom we show the trajectory of the baseline DEBiT-b agent on the same episode.
-In Fig.
-10 we show examples of our agent RANa-b navigating to two challenging ImageNav goals.
+In Fig.˜9 we show examples of our agent RANa-b static navigating to two challenging ImageNav goals.
 The figures visualize the observation (left), goal (highlighted in green), $N=8$ context elements, with the one selected at the given time step highlighted in red, and the agent path drawn on the scene floorplan on the right.
@@
 In this figure and the following ones, time is not sampled uniformly: we selected time steps, typically towards the end of the episode, where the agent makes interesting context selections to solve the episode.
-The episode in Fig.
-10 (left) is challenging because the environment is rather complex, the goal image is ambiguous and it is “hidden” in a terrace at the extreme perifery of the navigable space.
+The episode in Fig.˜9 (left) is challenging because the environment is rather complex, the goal image is ambiguous and it is “hidden” in a terrace at the extreme perifery of the navigable space.
 The baseline DEBiT-b agent, without context, does its best to explore the apartment and gets close to the terrace door where the goal is located, but does not manage to navigate to it.
@@
 Figure 10: Refer to caption  RANa-b successfully navigating to two goals in zero-shot Instance-ImageNav episodes - The agent observation is on the left, next to it is the original goal image from the Instance-ImageNav task (in cyan) and the retrieved goal image $\mathbf{g}^{\mathcal{R}}$ highlighted in green.
-The remaining data is organized and visualized as in Fig.
-10.
-Fig.
-11 shows RANa-b ImageNav agent successfully navigating in two zero-shot Instance-ImageNav episodes.
-Results are visualized in the same way as in Fig.
-10, with the exception of the goal image.
+The remaining data is organized and visualized as in Fig.˜9.
+Fig.˜10 shows RANa-b ImageNav agent successfully navigating in two zero-shot Instance-ImageNav episodes.
+Results are visualized in the same way as in Fig.˜9, with the exception of the goal image.
 In this case we have in fact two goal images, one from the original task (highlighted in cyan) which is taken with a sensor with different pose and intrinsics of the agent sensor, and the goal image $\mathbf{g}^{\mathcal{R}}$, highlighted in green, which is retrieved from the scene database $\mathcal{R}$.
@@
 Again we retrieve a goal image $\mathbf{g}^{\mathcal{R}}$ from $\mathcal{D}$ and the agent navigates in a 3D model of the environment.
-In Fig.
-12 we present an even more challenging scenario, where RANa navigates in two zero-shot Instance-ImageNav episodes, where the goal image is captured with a smartphone in a real environment.
-As mentioned in the main text, we have access to this environment and a 3D model of the scene is available.
+In Fig.˜11 we present an even more challenging scenario, where RANa navigates in two zero-shot Instance-ImageNav episodes, where the goal image is captured with a smartphone in a real environment.
+As mentioned in the main text, we assume having access to this environment and that a 3D model of the scene is available.
 The agent navigates to the retrieved goal image $\mathbf{g}^{\mathcal{R}}$ from the database $\mathcal{D}$ in the simulated reconstruction of the scene.
@@
 We can describe the behavior of our RANa agent in these cases as being “distracted” by context items visible by the agent at a given time step.
-Fig.
-13 shows a typical instance of this phenomenon: the baseline DEBiT agent (trajectory shown at the bottom) quickly gets to the goal, while RANa reaches the goal area, but starts turning on itself, cycling multiple times through relevant, but “distracting”, context images.
+Fig.˜12 shows a typical instance of this phenomenon: the baseline DEBiT agent (trajectory shown at the bottom) quickly gets to the goal, while RANa reaches the goal area, but starts turning on itself, cycling multiple times through relevant, but “distracting”, context images.